### v1

# Kareem Amin  Alex Bie  Weiwei Kong  Alexey Kurakin Natalia Ponomareva  Umar Syed  Andreas Terzis  Sergei Vassilvitskii Google

Abstract
We present an approach for generating differentially private synthetic text using large language models (LLMs), via private prediction. In the private prediction framework, we only require the output synthetic data to satisfy differential privacy guarantees. This is in contrast to approaches that train a generative model on potentially sensitive user-supplied source data and seek to ensure the model itself is safe to release. We prompt a pretrained LLM with source data, but ensure that next-token predictions are made with differential privacy guarantees. Previous work in this paradigm reported generating a small number of examples ($<10$) at reasonable privacy levels, an amount of data that is useful only for downstream in-context learning or prompting. In contrast, we make changes that allow us to generate thousands of high-quality synthetic data points, greatly expanding the set of potential applications. Our improvements come from an improved privacy analysis and a better private selection mechanism, which makes use of the equivalence between the softmax layer for sampling tokens in LLMs and the exponential mechanism. Furthermore, we introduce a novel use of public predictions via the sparse vector technique, in which we do not pay privacy costs for tokens that are predictable without sensitive data; we find this to be particularly effective for structured data.
1 Introduction
Differentially private mechanisms process a source dataset potentially containing sensitive user information and perform a useful computation — as simple as calculating a mean, or as complex as training an ML model — whose output can be safely shared while protecting the privacy of users who contributed to the dataset.
Perhaps the most general-purpose differentially private mechanism is one that produces a synthetic version of its input dataset, as the output of such a mechanism would be suitable for all the same purposes as the original dataset. For example, a private synthetic dataset can be used to train an ML model, but can also be used for auxiliary tasks such as feature engineering, hyperparameter tuning, and quality monitoring.
There has been recent interest in using large-language models (LLMs) to generate differentially private versions of text datasets. Existing approaches can be classified into several categories. Private fine-tuning methods privately adjust the parameters of an LLM on the input dataset, using an algorithm such as differentially private stochastic gradient descent (DP-SGD), and then prompt the LLM to generate similar text. Fine-tuning methods have been used to produce high-quality synthetic data, but the training procedure can be prohibitive, available only to those with the time, compute, and access necessary to train state-of-the-art LLMs containing billions of parameters.
Private prediction methods do not modify the LLM parameters at all. Instead, they directly prompt the LLM with text from the source dataset, asking for similar text in response, and then perturb the LLM’s token distribution (i.e., its last layer) to ensure that each sampled token, and thus the entire generated response, is private. Since no training is required, private prediction methods can quickly generate synthetic data, typically producing some data within minutes instead of hours, which allows for rapid prototyping and iteration. However, unlike private fine-tuning, the guarantees of private prediction methods degrade with the volume of data that is generated. Consequently, existing private prediction methods have mostly been used in applications that require only small amounts of synthetic data [Tang et al., 2024], sharply limiting their practicality.
In this paper we describe a new private prediction method that produces hundreds of times as much synthetic data as a state-of-the-art private prediction method, while maintaining a comparable privacy guarantee. Similar to some existing work, our method is based on running LLM inference on several subsets of the input data in parallel and privately aggregating their token distributions to generate synthetic text. However, our approach is distinguished by three novel algorithmic elements that lead to its improved performance:
  1. 1.
    Instead of protecting the privacy of the entire token distribution with Gaussian or Laplace noise, we leverage the uncertainty inherent in sampling to ensure privacy. We clip and aggregate token logits before standard softmax sampling — which is differentially private, since it can viewed as the exponential mechanism. Our approach induces much less distortion of the original token distributions to achieve the same level of privacy than prior work.
  2. 2.
    Previous work generated each token using a random subset of the input data, leveraging privacy amplification by subsampling in their analysis. This is computationally undesirable, as it requires repeated re-computation of the prefix for each decoding step, and limits scalability towards generating large synthetic corpora. Instead, we generate each synthetic example using a fixed disjoint subset of the input data, which yields substantial savings in privacy cost (by leveraging parallel composition) while allowing us to pay a linear amount of non-attention FLOPs, rather than quadratic, in terms of sequence length (via KV cache accelerated decoding).
  3. 3.
    Our method uses an auxillary token distribution from an LLM without access to sensitive data, and draws the next token from that distribution whenever it is very similar to the token distribution induced by the sensitive data. Our method incurs no privacy cost when outputting “obvious” tokens, and as a result, only a fraction of the tokens in the synthetic data are generated using sensitive data (as little as 20% in structured datasets). We leverage the sparse vector technique to privately calculate distributional similarity.
Taken together, the combination of these algorithmic techniques leads to significant improvements over prior work. Informally, (1) and (2) above keep our inference closely aligned to standard (non-DP) inference.
In our experiments, we generate private synthetic versions of publicly available, benchmark machine learning datasets, and then use the synthetic datasets for downstream classification and extraction tasks. Owing to the increased quantity and quality of our synthetic data, we improve over an existing state-of-the-art private prediction method in terms of downstream accuracy. Furthermore, while prior work in this paradigm only generated a small (<10) number of examples we demonstrate the ability to generate thousands of training examples, enough for fine-tuning downstream models.
Finally, since synthetic data is intended for a wide variety of applications, we also evaluate data quality using a metric that is orthogonal to performance on a downstream classification task. Specifically, we generate synthetic versions of a publicly available dataset containing highly structured data records, each of which is encoded as a JSON object. Our results demonstrate that the sparse vector technique helps preserve data structure at high privacy levels.
2 Related work
Private fine-tuning is widely used for synthetic text generation. Yue et al. [2023] created private synthetic versions of text datasets by using DP-SGD [Abadi et al., 2016] to fine-tune an LLM on the sensitive data. Kurakin et al. [2024] showed that parameter efficient approaches to fine-tuning, such as LoRA [Hu et al., 2022], can improve the quality of the synthetic data, since reducing the number of parameters also reduces the amount of noise injected into the optimization procedure. Wu et al. [2024a] took a two-stage approach: First they fine-tuned an LLM on a public dataset that closely resembled the sensitive data (which was itself generated by an LLM using carefully designed prompts), and then completed the fine-tuning process by running DP-SGD on the sensitive dataset. Concurrent to this work, Tran and Xiong [2024] describe a private fine-tuning approach for generating synthetic tabular data that is formatting compliant.
Private prediction [Dwork and Feldman, 2018] is an alternate approach to private machine learning that only guarantees the privacy of the predictions output by an ML model, and not the model itself. Private prediction has been applied to synthetic text generation by viewing each token sampled by an LLM as a ‘prediction’, and perturbing the LLM’s token distributions to ensure their privacy. Tang et al. [2024] added noise to several independent token distributions and averaged them, while Hong et al. [2024] selected the most popular token among the token distributions using the LimitedDomain mechanism [Durfee and Rogers, 2019]. These methods can avoid the time, compute, and access required to fine-tune an LLM with billions of parameters. However, a privacy loss is suffered for each token produced in this manner. As a result, previous work has only been able to generate a very small number of synthetic examples at reasonable privacy levels (fewer than 10). Other work has applied private prediction techniques to LLMs [Majmudar et al., 2022, Duan et al., 2023], including in combination with fine-tuning [Ginart et al., 2022, Flemings et al., 2024], but not for the purpose of synthetic text generation.
Finally, another distinct set of approaches are private filtering methods. Private filtering methods operate directly on whole LLM responses and a large corpus of public data that does not require protection. Yu et al. [2024] and Xie et al. [2024] used the sensitive responses to privately select similar responses from the public dataset. Similarly, Wu et al. [2024b] aggregate response embeddings and select the public response that is closest in embedding space. [^1] Wu et al. [2024b] also proposed a non-filtering approach based on privately selecting common keywords among the sensitive data and using them to prompt an LLM. One limitation of filtering methods is that the menu of possible responses is constructed without signal from the new source dataset.
3 Method
Before describing our algorithm for generating private synthetic text, we review the standard algorithm for LLM inference. Let $\mathcal{X}$ be the token vocabulary (i.e., the set of all possible tokens), and let $v=|\mathcal{X}|$ be the vocabulary size. A token sequence is an element of $\mathcal{X}^{*}$, and a logit vector is an element of $\mathbb{R}^{v}$ (one logit per token in the vocabulary). If $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$ are token sequences then we write $\mathbf{x}_{1}\mathbf{x}_{2}\in\mathcal{X}^{*}$ to denote their concatenation.
Standard LLM inference.
[…]
Our algorithm.
One straightforward approach to generating a synthetic version of a sensitive piece of text would be to prompt an LLM with ‘Please generate text similar to: <sensitive text>’. However, this could easily lead to a privacy violation, as the response could retain the semantics of the input sensitive text.
Figure 1: Visualization of Algorithm 1 for a single token in a single batch.
Algorithm 1 describes our method for privately generating a dataset of synthetic examples $X$ from a dataset of sensitive prompts $D$. Each prompt in $D$ resembles the sample prompt given above. But instead of using a single prompt to generate a synthetic example, the algorithm uses a batch of the prompts to run several LLM inferences in parallel. Each synthetic example is generated one token at a time, with the average of the logit vectors across the inferences defining the distribution from which the next token is randomly selected. Before averaging, the logits are clipped and re-centered using the function
$\operatorname{clip}_{c}(\mathbf{z})_{i}=\max\left(-c,\mathbf{z}_{i}-\max_{j}{\{\mathbf{z}_{j}\}}+c\right)$ (1)
which maps each component of $\mathbf{z}$ into $[-c,c]$. Forcing each logit to lie in a bounded range is key to proving the privacy guarantee for our algorithm (see §4). While several functions can achieve this purpose, Eq. (1) has an additional desirable property: If the components of $\mathbf{z}$ can be shifted by a constant so that they all lie in the interval $[-c,c]$, then $\operatorname{clip}_{c}(\mathbf{z})$ is one such shift. This property is desirable because the distribution $\operatorname{softmax}(\mathbf{z})$ is invariant to any constant shift of $\mathbf{z}$. We also found that Eq. (1) performed better empirically than other functions we considered. For example, regular clipping to the range $[-c,c]$ without recentering requires twice as large $c$ to sample without distortion (see Appendix B).
Algorithm 1 Generate private synthetic examples using an LLM
1:Parameters: LLM logit function $\operatorname{logits}(\cdot)$, public prompt $\mathbf{p}_{\operatorname{public}}$, expected batch size $s$, maximum number of private tokens per batch $r$, clipping value $c$, noise level $\sigma$, distance function $d$, threshold $\theta$, public temperature $\tau_{\operatorname{public}}$, private temperature $\tau$
2:Input: Dataset of sensitive prompts $D$
3:Output: Dataset of synthetic examples $X$
4:Let $\mathcal{S}$ be a partition of $D$ into disjoint batches
5:for each batch $S\in\mathcal{S}$ do
6:     $\hat{\theta}\leftarrow\theta+\textrm{Laplace}(\sigma)$
7:     $t\leftarrow 0$ private token counter
8:     while $t<r$ do
[…]
10:         while $\mathbf{x}$ does not end with <eos> do
11:              $Z\leftarrow\{\operatorname{logits}(\mathbf{p}\mathbf{x}):\mathbf{p}\in S\}$
12:              $\mathbf{z}_{\operatorname{public}}\leftarrow\operatorname{logits}(\mathbf{p}_{\operatorname{public}}\mathbf{x})$
13:              $\hat{d}\leftarrow d(Z,\mathbf{z}_{\operatorname{public}})+\textrm{Laplace}(2\sigma)$
14:              if $\hat{d}>\hat{\theta}$ then
15:                  $\bar{\mathbf{z}}\leftarrow\frac{1}{s}\sum_{\mathbf{z}\in Z}\operatorname{clip}_{c}(\mathbf{z})$
16:                  $x\sim\operatorname{softmax}(\bar{\mathbf{z}}/\tau)$
17:                  $\hat{\theta}\leftarrow\theta+\textrm{Laplace}(\sigma)$
18:                  $t\leftarrow t+1$
19:              else
20:                  $x\sim\operatorname{softmax}(\mathbf{z}_{\operatorname{public}}/\tau_{\operatorname{public}})$
21:              Append $x$ to $\mathbf{x}$
22:         Add $\mathbf{x}$ to $X$
23:Return $X$.
Since the average logit vector is computed using a set of sensitive prompts, each token selected from a distribution determined by the average logit vector incurs a privacy cost. To minimize this cost, Algorithm 1 also has access to a non-sensitive public prompt, $\mathbf{p}_{\operatorname{public}}$, and uses this prompt to generate the next token whenever doing so does not significantly change the distribution from which the next token is drawn. The distance function used to make this determination is
$d(Z,\mathbf{z}_{\operatorname{public}})=\left\lVert\frac{1}{s}\sum_{\mathbf{z}\in Z}p_{\mathbf{z}}-p_{\mathbf{z}_{\operatorname{public}}}\right\rVert_{1}$ (2)
where $p_{\mathbf{z}}=\operatorname{softmax}(\mathbf{z})$ is the token distribution corresponding to logit vector $\mathbf{z}$. When this distance is small, Algorithm 1 outputs a public token instead of a private token. The privacy guarantee for Algorithm 1 leverages the analysis of the sparse vector technique [Dwork et al., 2009], and shows that while privacy degrades with the number of private output tokens, it is independent of the number of public output tokens (see §4). Empirically, we observe that the fraction of output tokens that must be private in order to generate high-quality synthetic data can be only 20% for some datasets.
Note that the first step of Algorithm 1 partitions the input dataset of sensitive prompts into disjoint batches. We do not prescribe a procedure for assigning prompts to batches in Algorithm 1 since many batching approaches are admissible as long as they satisfy a minor technical assumption required for the privacy analysis of Algorithm 1, which we explain in §4. While the batches are not required to be any particular size, the algorithm makes most efficient use of the data if each batch has size equal to the expected batch size $s$. And while prompts can be batched together (almost) arbitrarily, more tailored batching can lead to better synthetic data quality. For example, in the experiments in §5, where we generate synthetic versions of ML training datasets, each sensitive prompt contains a label. In those experiments we assign prompts with the same label to the same batch.
Relationship to prior work.
Two major features of Algorithm 1 are that it leverages the inherent randomness of token sampling to guarantee privacy, and that it further reduces privacy cost by using public data to generate a portion of the synthetic data. Some prior work also incorporated these algorithmic ideas, but with key differences. Instead of clipping logits to ensure that the token sampling is private, Majmudar et al. [2022] mixed each sensitive token distribution with the uniform distribution. This approach induced a dependence on the vocabulary size in their privacy guarantee, and since LLM vocabularies are typically very large, the resulting privacy guarantee was quite weak: Majmudar et al. [2022] noted that setting the differential privacy parameter $\varepsilon$ (see Definition 1) lower than 50 produced synthetic data that was “unusable”. Flemings et al. [2024] guaranteed the privacy of token sampling by mixing each sensitive token distribution with a public token distribution, but their approach was based on aggregating a set of fine-tuned models, not a set of prompts. Neither Majmudar et al. [2022] nor Flemings et al. [2024] had a goal of generating synthetic data.
Tang et al. [2024] found that limiting the token vocabulary to a fixed set of the most popular 100 public tokens caused their synthetic data generation algorithm to exhibit greater stability. However, if the sensitive data contains many tokens that are rare in public data, their approach cannot produce synthetic data that is very similar to the sensitive data. By contrast, our approach compares public and private token distributions on-the-fly, and determines which one to use for sampling the next token by balancing a trade-off between privacy and quality. Also, Tang et al. [2024] used a different random subset of prompts to generate each token, and left as an open problem how to use a single subset to generate every token in a synthetic example. Our algorithm resolves this open problem, and consequently yields both improved privacy and greater computational efficiency (see §6).
4 Privacy analysis
In this section we prove that Algorithm 1 preserves the privacy of the sensitive prompts it uses to generate synthetic examples.
Let $\mathcal{D}$ be the set of all possible prompt datasets. A mechanism is a randomized function with domain $\mathcal{D}$. Note that Algorithm 1 is a mechanism. We say that a pair of prompt datasets $D,D^{\prime}\in\mathcal{D}$ are neighbors if there exists a prompt $\mathbf{p}$ such that $D=D^{\prime}\cup\{\mathbf{p}\}$ or $D^{\prime}=D\cup\{\mathbf{p}\}$. In the differential privacy literature this is commonly referred to as the add/remove neighbor relation.
Definition 1 (Dwork et al. [2006]).
A mechanism $M$ satisfies $(\varepsilon,\delta)$-differential privacy if $\Pr[M(D)\in O]\leq e^{\varepsilon}\Pr[M(D^{\prime})\in O]+\delta$ for any neighboring datasets $D,D^{\prime}\in\mathcal{D}$ and subset $O$ of the range of $M$.
Theorem 1 below provides a differential privacy guarantee for Algorithm 1. The proof of Theorem 1 requires a technical assumption about how the prompts are partitioned into batches in the first step of the algorithm.
Assumption 1.
In Algorithm 1, the assignment of a prompt to a batch depends only on the prompt itself, and not on the other prompts.
The most straightforward way to satisfy Assumption 1 is to apply a hash function to each prompt and then use the hash value to determine its assigned batch. For example, if $h$ is the hash value, $n$ is the number of prompts and $s$ is the expected batch size, then we can assign the prompt to the $(h\mod\frac{n}{s})$th batch. If we want to batch together prompts that share a certain attribute (like a label), we can apply another hash function to that attribute and concatenate the hash values. Using hash functions for batch assignment can lead to batches whose sizes differ from the expected batch size $s$, but this does not impact the validity of Theorem 1.
Theorem 1 (Privacy of Algorithm 1).
Suppose Assumption 1 holds. Let $\rho=r\left(\frac{1}{2}\left(\frac{c}{s\tau}\right)^{2}+\frac{2}{(s\sigma)^{2}}\right)$. For all $\varepsilon\geq 0$, Algorithm 1 satisfies $(\varepsilon,\delta)$-differential privacy, where
$\delta=\inf_{\alpha\in(1,\infty)}\frac{e^{(\alpha-1)(\alpha\rho-\varepsilon)}}{\alpha-1}\left(1-\frac{1}{\alpha}\right)^{\alpha}.$
Also, for all $\delta\in(0,1]$, Algorithm 1 satisfies $(\varepsilon,\delta)$-differential privacy, where
$\varepsilon=\rho+\sqrt{4\rho\log(1/\delta)}.$
The proof is in Appendix C and makes use of sharp privacy analyses of: (1) zCDP to approximate DP conversion [Canonne et al., 2020]; and (2) zCDP bounds for the exponential mechanism [Cesar and Rogers, 2021].
5 Experiments
Gemma 1.1 2B IT [Team, 2024] is the data generator in all private prediction experiments. We choose it due to its lightweight, open-source JAX implementation that makes easy to implement and share sampling algorithms. [^2] https://github.com/google-deepmind/gemma Tables 1(a) and 1(b) give an overview of datasets and models used.
Dataset $n_{\text{train}}$ Description
AGNews 120,000 4-way news classification
TREC 5452 6-way query classification
DBPedia 560,000 14-way topic classification
MIT-G 2,953 Movie genre extraction
MIT-D 1,561 Movie director extraction
IMDB 25,000 2-way review classification
Yelp 560,000 2-way review classification
WikiMoviesJSON 27,412 JSON with 6 fields
(a) Overview of datasets used.
Model Usage
Gemma 2B 1.1 IT Generation; private prediction
LaMDA 8B Generation; DP fine-tuning
GPT-3 babbage-002 Evaluation; in-context learning
BERT-Base 12/768 110M Evaluation; fine-tuning
(b) Overview of models used.
Table 1: Overview of datasets and models used. Datasets are benchmark classification and extraction tasks used in prior work on private synthetic text generation, with the exception of WikiMoviesJSON, which is used for structured data experiments. LaMDA and Gemma are used for synthetic data generation, while the other models are used to measure how useful our synthetic data is for improving accuracy on real test data.
We perform 3 sets of experiments, targeting various datasets and utility criteria:
  • In-context learning (§5.1); We generate examples to use as in-context exemplars for prompting an LLM. We report downstream accuracy on real test examples, when prompted with synthetic data, on 3 classification tasks (AGNews [Zhang et al., 2015], DBPedia [Zhang et al., 2015], TREC [Voorhees and Tice, 2000]) and 2 extraction tasks (MIT-G, MIT-D [Liu et al., 2012]).
  • Fine-tuning (§5.2); We generate synthetic examples to use for fine-tuning a BERT classifier. We report downstream accuracy on real test examples for 3 classification tasks (IMDB [Maas et al., 2011], Yelp [Zhang et al., 2015], AGNews [Zhang et al., 2015]).
  • Structured data (§5.3); We generate examples that must adhere to structural constraints to be useful synthetic data. We consider a JSON generation task (WikiMoviesJSON [Rust, 2024]), evaluating structure preservation.
5.1 In-context learning
Experimental setup.
Using our method, we generate 90-1500 examples using Gemma 2B 1.1 IT. We compare against real examples, and results reported in the prior work of Tang et al. [2024], where they generated 4-shot examples for in-context learning. [^3] It is no longer possible to reproduce their results, due to changes in the OpenAI API since publication: GPT-3 babbage is now deprecated, and it is no longer possible to query for top 100 logprobs, which is required by their method. To evaluate generated synthetic data, we put synthetic examples in the context window before querying with the real test example, as shown in Figure 2.
1Classify the following examples: 2Text: lorem ipsum # synthetic text 1 3Answer: label 4... 5Text: sed do eiusmod # synthetic text n 6Answer: label 7 8Text: excepteur sint # test text 9Answer:
Figure 2: Example of $n$-shot in-context learning evaluation for synthetic data.
We perform this evaluation with GPT-3 babbage-002 which has a 16K context window. We report results on AGNews, DBPedia, TREC, MIT-G, and MIT-D using the implementation of Zhao et al. [2021]. [^4] https://github.com/tonyzhaozh/few-shot-learning. Following the work of Tang et al. [2024], we enable contextual calibration [Zhao et al., 2021] for classification but not extraction tasks. Our evaluation setup is a best-effort reproduction of their setup, which is no longer possible to completely reproduce due to changes to OpenAI API access (see Table 2 caption for more details). Due to cost, we follow prior work [Bertsch et al., 2024, Ratner et al., 2023, Lu et al., 2022, Zhao et al., 2021] and opt to subsample test sets to 250 test examples. We run 3 seeds of label-stratified sampling of exemplars from synthetic/real data.
GPT-3 babbage-002 Acc. (%)*
$\varepsilon$ Method Shots Reported in Model AGNews DBPedia TREC MIT-G MIT-D
0 Zero shot 0 This work - $24.8_{0.0}$ $12.0_{0.0}$ $28.4_{0.0}$ $29.6_{0.0}$ $28.8_{0.0}$
$\infty$ Real data 4 This work - $75.3_{3.0}$ $73.6_{0.3}$ $34.9_{5.0}$ $56.0_{2.0}$ $83.1_{5.3}$
64 $84.7_{1.5}$ $92.5_{1.6}$ $50.3_{6.1}$ $56.4_{5.4}$ $89.1_{0.7}$
Tang et al. [2024] 4 Tang et al. [2024]* GPT-3 babbage $69.3_{4.8}$ $82.3_{3.7}$ $50.6_{6.9}$ $54.4_{7.0}$
Ours 4 This work Gemma 1.1 2B IT $76.8_{4.8}$ $72.3_{2.5}$ $38.8_{6.0}$ $47.7_{2.5}$ $81.7_{2.4}$
64 $77.5_{1.8}$ $91.5_{1.7}$ $57.9_{3.4}$ $56.4_{1.2}$ $87.1_{0.2}$
$1$ Tang et al. [2024] 4 Tang et al. [2024]* GPT-3 babbage $64.1_{3.9}$ $81.2_{3.0}$ $50.7_{4.1}$ $46.3_{7.8}$ $69.2_{7.9}$
Ours 4 This work Gemma 1.1 2B IT $75.9_{3.5}$ $75.1_{0.5}$ $39.2_{3.7}$ $47.1_{6.0}$ $84.5_{1.0}$
64 $78.7_{1.8}$ $90.4_{2.6}$ $53.6_{1.3}$ $51.6_{2.3}$ $86.4_{0.6}$
Table 2: In-context learning results with GPT-3 babbage-002. We report mean and standard deviation over 3 random samplings (equally many from each label for classification; fully random for extraction) of synthetic/real data. (*) Note: For the results reported in Tang et al. [2024], they use GPT-3 babbage (now deprecated; we use GPT-3 babbage-002) as the in-context learner, and use the top 100 logprobs for contextual calibration (only top 5 are available now); while not directly comparable, we report their results for context.
Results.
Results are presented in Table 2. Our gains in quantity while maintaining quality are realized in terms of 64-shot in-context learning accuracy. In some cases, we can generate more examples, but we limit ourselves to 64 for these evaluations for cost and efficiency reasons. Our results at 64 shots are comparable to real data at 64 shots. Notably, our synthetic data at 64 shots improves over real data at 4 shots – which is roughly an upper bound on the performance of methods limited to generating 4 examples (e.g., Tang et al. [2024]). We also improve over results reported in Tang et al. [2024], although we note that there are differences in the experimental setup.
5.2 Fine-tuning
We achieve significant improvements over the best available private inference method for in-context learning tasks. Since our method is capable of generating thousands of synthetic examples at reasonable privacy budgets, it is natural to ask whether it can compete with state-of-the-art private fine-tuning methods, which can generate infinitely many synthetic examples once the up-front costs of model training are paid. This makes them capable of producing enough data to train downstream classification models.
Experiment setup.
We use our approach to generate a large quantity of synthetic data for the purposes of fine-tuning 110M BERT-Base models. We consider 3 classification tasks used in prior work on private fine-tuning [Kurakin et al., 2024]), following the exact same evaluation procedure. We omit comparison to prior private prediction work (e.g. [Tang et al., 2024]), as they only generate 4 examples which is insufficient for fine-tuning.
Results.
Main results are presented in Table 3. Across various datasets and privacy levels, we generate between 2.5K (IMDB, $\varepsilon$=1) and 200K (Yelp, $\varepsilon$=10) examples for fine-tuning. Prior work generating fewer than $10$ examples using private prediction were unable to compare with private fine-tuning on these tasks at all. While there remains a gap between the best fine-tuning and best private inference methods on downstream classification tasks, we achieve reasonable performance, even out-performing or matching the baseline of privately tuning all the parameters in the model reported in Kurakin et al. [2024].
BERT Acc. (%)
IMDB @ $\varepsilon$ Yelp @ $\varepsilon$ AGNews @ $\varepsilon$
Method Reported in Model $\infty$ $1$ $3$ $10$infty $1$ $3$ $10$infty $1$ $3$ $10$
Real data [Kurakin et al., 2024] - $93.7_{0.1}$ $97.6_{0.1}$ $93.7_{0.1}$
Fine-tune [Kurakin et al., 2024] LaMDA 8B $93.2_{0.2}$ $79.1_{1.7}$ $83.9_{0.6}$ $84.0_{0.7}$ $95.9_{0.1}$ $84.1_{0.3}$ $84.6_{0.1}$ $84.2_{0.3}$ $91.1_{0.1}$ $65.7_{2.9}$ $65.3_{2.7}$ $65.1_{5.3}$
Prompt-tune $92.0_{0.1}$ $88.1_{0.4}$ $87.4_{0.2}$ $90.7_{0.2}$ $93.9_{0.1}$ $94.1_{0.1}$ $93.5_{0.1}$ $94.1_{0.1}$ $88.3_{0.3}$ $83.9_{0.8}$ $86.2_{0.2}$ $86.9_{0.1}$
LoRA $91.6_{0.2}$ $90.0_{0.3}$ $90.6_{0.2}$ $91.3_{0.2}$ $96.4_{0.1}$ $95.5_{0.1}$ $95.6_{0.1}$ $95.9_{0.1}$ $91.8_{0.2}$ $89.4_{0.1}$ $89.6_{0.1}$ $90.0_{0.1}$
Ours This work Gemma 1.1 2B IT $83.6_{2.9}$ $82.7_{2.1}$ $83.6_{1.9}$ $85.5_{2.3}$ $91.8_{0.6}$ $91.1_{0.2}$ $91.6_{0.8}$ $92.6_{0.2}$ $81.2_{1.2}$ $79.8_{1.8}$ $79.3_{2.1}$ $79.8_{0.3}$
+ SVT This work Gemma 1.1 2B IT - $84.3_{1.1}$ $84.4_{1.5}$ $85.0_{1.0}$ $88.4_{0.6}$ $89.1_{0.3}$ $89.0_{1.9}$ $79.2_{0.3}$ $79.8_{0.4}$ $80.4_{0.6}$
Table 3: Results of fine-tuning on real and synthetic data with BERT. We report mean and standard deviation over 3 runs of downstream fine-tuning and evaluation. We compare to results reported in [Kurakin et al., 2024] that fine-tunes a synthetic data generator with DP-SGD. We generate 2.5-200K examples with private prediction, which suffices to train reasonably performing models on.
Limited data regime.
We additionally consider the limited data regime. In Appendix A we present experiments on AGNews1K, a 1024-subsample of AGNews. Our method, which employs parallel composition, is “pay-as-you-go”, i.e., we can put in a small amount of data to get out a small amount, while preserving quality. On the other hand, fine-tuning based approaches necessarily pay upfront to ensure the model and all future generations are private. This means that without sufficient data, all outputs will be low quality. Results in Table 5 demonstrate that our private prediction method generates more useful examples for in-context learning in this regime.
5.3 Structured data
We conclude our experiments with a demonstration of the lift in performance provided by using the sparse vector technique (SVT) against a public prompt. Informally, the privacy loss of our method only scales with the information density of a new example vis-a-vis the public prompt. This contrasts with other private inference methods that incur privacy loss on every token. This is especially useful for structured data, where we avoid incurring privacy loss on syntactic elements of the data.
Experiment setup.
For JSON generation, we evaluate on a dataset of information about American movies scraped from Wikipedia [Rust, 2024]. Entries contain fields such as title, year, cast, and extract (a short synopsis). We lightly curate the data: we omit uninteresting fields (i.e., thumbnail dimensions and URLs) and remove entries with incomplete entries. We refer to the resulting 34,266 JSON examples with 6 fields as WikiMoviesJSON. We evaluate two criteria: the rate at which output generated constitutes well-formed JSON (Parses (%)), and rate at which the output passes basic schema validation (Validates (%)). This includes checks such as: no extra fields, all required fields are present, values are the correct type, and other custom constraints (e.g. no whitespace in the href field).
Results.
Results are in Table 4. Targeting a large number of examples at small $\varepsilon$ necessitates increases in the sampling temperature $\tau$, to ensure privacy, but compromises the well-formed-ness of outputs. For structured generation, there is a large amount of tokens that (a) are crucial to get right for structure preservation, and (b) easily predictable without access to sensitive data. Here the SVT enables us to get these tokens reliably and for free, leading to better generation quantity.
$\varepsilon$ Method $\tau$ Parses (%) Validates (%) $m$
1 Ours 2 $80.6_{1.3}$ $74.2_{1.9}$ $94.3_{1.2}$
2.5 $4.9_{1.1}$ $1.5_{0.1}$ $138.0_{7.5}$
+ SVT, $\theta$ = 0.9 2 $91.7_{2.1}$ $88.6_{3.2}$ $289.7_{19.4}$
2.5 $74.1_{2.7}$ $64.0_{4.1}$ $356.7_{25.9}$
+ SVT, $\theta$ = 1.5 2 $95.5_{1.0}$ $93.1_{0.7}$ $893.0_{20.2}$
2.5 $79.3_{1.0}$ $72.7_{1.4}$ $1178.3_{10.1}$
Table 4: Results for generating JSON records from WikiMoviesJSON. We report mean and standard deviation over 3 runs of dataset generation. $\tau$ refers to the sampling temperature, and $m$ refers to the number of raw samples produced (before parsing and validation checks). The batch size used is 255. We present results at two different SVT thresholds $\theta$, and see gains in structure preservation and quantity.
6 Discussion
We believe that our significantly improved performance relative to Tang et al. [2024] is primarily attributable to two algorithmic innovations.
First, for each generated token, Tang et al. [2024] preserve the privacy of the entire distributions from which the token is sampled (by taking argmax), even though only the token itself is included in the synthetic data. By contrast, our method uses a discrete choosing mechanism, the exponential mechanism. As a result, we do not need to maintain a DP version of the entire token distribution to release a single token. This decision leads to significantly lower noise requirements, as a straightforward calculation reveals. Empirically, we obtained good synthetic data quality with $s=250$, $\tau=2$, $c=10$ and $\delta=10^{-6}$. In order to switch to the Gaussian mechanism using its standard $(\varepsilon,\delta)$-DP guarantee, and achieve comparable privacy guarantees we would would require $\sigma\approx 0.53$ to achieve a comparable privacy guarantee. (See Appendix D). Better analyses of the Gaussian mechanism exist, but do not offer much help. Using the improved analysis in Balle and Wang [2018] to attain the same $\varepsilon$ would require $\sigma\approx 0.34$. Conducting the analysis so that both mechanisms have equivalent privacy loss under zCDP yields $\sigma=0.2$. These are all very large noise magnitudes relative to probabilities in $[0,1]$. [^5] To put independent noise of magnitude $\sigma=0.2$ into perspective: suppose the ground truth next-token prediction is deterministic, i.e., $\bar{\mathbf{p}}=[1,0,...,0]\in\mathbb{R}^{v}$, $v$ = 256128 in the case of Gemma. Now with probability $\geq 0.15$, the noised distribution ${\widetilde{\mathbf{p}}}$ has ${\widetilde{\mathbf{p}}}_{1}<0.8$. Each other $\mathbf{p}_{i}$ is $\geq 0.8$ w.p. $\geq 3\cdot 10^{-5}$ independently. Hence the probability of one of these being promoted to argmax is $\geq 0.15\cdot(1-(1-3\cdot 10^{-5})^{v-1})\approx 0.15$. At this rate, the chance of generating a 30 token span without a corruption is $<1\%$.
Secondly, Tang et al. [2024] generated each token using a different random sample of the sensitive prompts, which is computationally very expensive, as it prevents the use of KV cache-accelerated decoding, since the cache is invalidated upon every resample. While resampling less often would be more practical, Tang et al. [2024] noted that in this case the privacy amplification benefits of subsampling would not be adequately realized, and characterized this limitation as the “main weakness” of their approach. Instead, our method generates each synthetic example using a fixed disjoint subset of the sensitive prompts, allowing us to leverage parallel composition in our analysis, and thus avoid this privacy versus computation tradeoff.
7 Conclusion
As proprietary models become increasingly powerful, we anticipate more practitioners will be able to generate inferences from state-of-the-art models, while fewer practitioners will be able to train networks that perform like state-of-the-art models. This makes it increasingly important to develop private prediction methods that compete with private fine-tuning.
We demonstrate that private prediction can be used to generate large amounts of synthetic text with reasonable differential privacy guarantees. We produce 2-3 orders of magnitude more private synthetic data than what was demonstrated in prior work in this paradigm. Access to more synthetic data lets us fine-tune downstream models, as well as yields performance improvements via many-shot in-context learning. Furthermore, we introduce a novel use of public models in which we are able to sample predictable tokens at no privacy cost, which is particularly effective for structured data.
Limitations
While our work demonstrates that private prediction is a practical technique for privately generating a large volume of high-quality synthetic data, there remains a small gap between our results and the results obtained from privately fine-tuning the parameters of the LLM. Currently, private prediction methods pay a privacy cost for every generated token, while private fine-tuning methods do not. We view correcting this limitation as a very important open problem. Finally, any method for ensuring data privacy will inevitably entail some loss of data utility.
Author contributions
  • Alex B is the main contributor. He implemented the method, tested variants to optimize utility and privacy, and ran most of the experiments. He also proposed the use of sparse vector.
  • Umar proposed the method, the use of sampling to preserve privacy, and conducted the theoretical analysis.
  • Umar and Kareem framed the structure of the paper and led writing.
  • Kareem proposed parallel composition. He also assisted with the privacy analysis.
  • Natalia proposed logits recentering.
  • Weiwei and Alexey provided infrastructure support and code for running experiments. Alexey suggested the limited data experiments and ran the fine-tuning baselines.
  • Natalia, Andreas, and Sergei advised the project.
  • Everyone contributed to discussing, interpreting, and iterating on experiment results as well as project management.
Appendix A Private prediction beats fine-tuning in the limited data regime
We do LoRA fine-tuning with DP-SGD on AGNews1K, with the same setup that beats our method in the full data regime. We sample synthetic data from the fine-tuned model. We also run our private prediction method on AGNews1K. We evaluate performance on 4 and 16 shot in-context learning with GPT-3 babbage-002 (the same experimental setting as §5.1).
$\varepsilon$ Method Shots Model Acc. (%)
1 LoRA 4 LaMDA 8B $63.3_{8.0}$
16 $68.1_{5.9}$
Ours 4 Gemma 2B 1.1 IT $73.9_{8.3}$
16 $80.1_{2.5}$
Table 5: Results on AGNews1K, a 1024-subsample of AGNews. Our method is “pay-as-you-go”, and is capable of generating a few high quality examples for in-context learning in this regime. On the other hand fine-tuning does worse due to the stricter requirement that all future model outputs must be private. We run 16 shot since we only generate 38 examples/since 16 examples fills up the context length for LoRA.
Appendix B Design choices
B.1 Logits clipping function
In Figure 3, we compare results for different logits clipping functions. The baseline approach it to clip all logits to the interval $[-c,c]$ before aggregation and softmax – we refer to this as “fixed interval clipping”. Alternatively, we can clip to the range $[\max_{j}\{\mathbf{z}_{j}\}-2c,\max_{j}\{\mathbf{z}_{j}\}]$ and then translate to the interval $[-c,c]$ (Equation 1). In Figure 3 we plot the distortion as a consequence of clipping in terms of L1 error, and find that the latter approach allows us clip more than twice as aggressively, thus improving the privacy guarantee, without compromising utility.
(a) Distribution of L1 error induced by fixed interval clipping.
(b) Distribution of L1 error induced by clipping with recentering.
Figure 3: We sample a few hundred tokens using logits aggregation with no clipping. At each sampling step, we compute the L1 distances between the post-softmax distributions of aggregated clipped logits vs. aggregated unclipped logits, at various settings of $c$, and plot them in a histrogram. We observe less error, at lower choices of $c$ when clipping with recentering (note the $x$-axis scales).
Appendix C Proof of Theorem 1
Our proof of Theorem 1 is organized into sections. §C.1 provides basic definitions. §C.2 and §C.3 establish key results related to composition and sensitivity. §C.4 proves the privacy of simpler mechanisms that each account for a portion of the functionality of Algorithm 1. C.5 puts all the pieces together and completes the proof.
C.1 Definitions
In §4 we defined neighboring prompt datasets. We extend the definition to arbitrary sets.
Definition 2.
Let $\mathcal{U}$ be a set. Let $S,S^{\prime}\subseteq\mathcal{U}$. We say that $S$ and $S^{\prime}$ are neighbors if there exists $u\in\mathcal{U}$ such that $S=S^{\prime}\cup\{u\}$ or $S^{\prime}=S\cup\{u\}$.
The sensitivity of a function is an upper bound on how much its value can change over neighbors.
Definition 3.
Let $\mathcal{U}$ be a set. Let $k\geq 1$. Let $f:2^{\mathcal{U}}\rightarrow\mathbb{R}^{k}$. The sensitivity of $f$ is
$\sup_{S,S^{\prime}}\left\lVert f(S)-f(S^{\prime})\right\rVert_{\infty}$
where the supremum is over neighbors $S,S^{\prime}\in\mathcal{U}$.
Zero-concentrated differential privacy (zCDP) is a relaxation of $\varepsilon$-differential privacy.
Definition 4 (Bun and Steinke [2016]).
A mechanism $M$ satisfies $\rho$-zCDP if
$D_{\alpha}(M(D)\parallel M(D^{\prime}))\leq\rho\alpha$
for all $\alpha>1$ and neighboring datasets $D,D^{\prime}\in\mathcal{D}$, where $D_{\alpha}(P\parallel Q)$ is Rényi divergence of order $\alpha$ betweeen distributions $P$ and $Q$.
C.2 Composition
Zero-concentrated differential privacy obeys a simple sequential composition rule.
Lemma 1.
If mechanisms $M_{1}$ and $M_{2}$ satisfy $\rho_{1}$-zCDP and $\rho_{2}$-zCDP, respectively, then the sequential composition of $M_{1}$ and $M_{2}$ satisfies $(\rho_{1}+\rho_{2})$-zCDP.
Parallel composition is a well-known technique in differential privacy that is useful for establishing privacy guarantees in scenarios where a mechanism is independently applied to disjoint subsets of a dataset. Many versions of parallel composition require that the subsets are chosen in a fully data-independent manner. We show that the same result holds under a weaker assumption.
Lemma 2.
Let $k$ be a positive integer. Let $f$ be a function that maps prompts into $[k]$. For any dataset of prompts $D$ and $i\in[k]$ let
$D_{i}=\{\mathbf{p}\in D:f(\mathbf{p})=i\}.$
Let $M$ be a mechanism that satisfies $\rho$-zCDP. If $\widehat{M}$ is the mechanism defined by
$\widehat{M}(D)=(M(D_{1}),\ldots,M(D_{k}))$
then $\widehat{M}$ satisfies $\rho$-zCDP.
Proof.
Let $D,D^{\prime}\in\mathcal{D}$ be neighboring datasets. Without loss of generality assume $D=D^{\prime}\cup\{\mathbf{p}\}$, where $\mathbf{p}$ is a prompt. There exists $j\in[k]$ such that $D_{i}=D^{\prime}_{i}$ for all $i\neq j$ and $D_{j}=D^{\prime}_{j}\cup\{\mathbf{p}\}$. We have for all $\alpha>1$
$D_{\alpha}(\widehat{M}(D)\parallel\widehat{M}(D^{\prime}))$ $=\sum_{i=1}^{k}D_{\alpha}(M(D_{i})\parallel M(D^{\prime}_{i}))$
$=D_{\alpha}(M(D_{j})\parallel M(D^{\prime}_{j}))$
$\leq\rho\alpha\qed$
C.3 Sensitivity analysis
In this we compute the sensitivity of several functions used in Algorithm 1. Each function depends on a set of logit vectors. Recall that a logit vector is an element of $\mathbb{R}^{v}$. Let
$\ell(Z)=\frac{1}{s}\sum_{\mathbf{z}\in Z}\operatorname{clip}_{c}(\mathbf{z})$
where $\operatorname{clip}_{c}(\cdot)$ was defined in Eq. (1). Also recall the distance function defined in Eq. (2):
$d(Z,\mathbf{z})=\left\lVert\frac{1}{s}\sum_{\mathbf{z}^{\prime}\in Z}p_{\mathbf{z}^{\prime}}-p_{\mathbf{z}}\right\rVert_{1}$
where $p_{\mathbf{z}}=\operatorname{softmax}(\mathbf{z})$.
Lemma 3.
The function $\ell$ has sensitivity $\frac{c}{s}$, and for all $\mathbf{z}\in\mathbb{R}^{v}$, the function $d(\cdot,\mathbf{z})$ has sensitivity $\frac{1}{s}$.
Proof.
Let $Z,Z^{\prime}\subseteq\mathbb{R}^{v}$ be neighbors. Let $\mathbf{\tilde{z}}\in\mathbb{R}^{v}$ be the logit vector they do not have in common. We have
$\left\lVert\ell(Z)-\ell(Z^{\prime})\right\rVert_{\infty}=\frac{1}{s}\left\lVert\operatorname{clip}_{c}(\mathbf{\tilde{z}})\right\rVert_{\infty}\leq\frac{c}{s}.$
We also have
$\left|d(Z,\mathbf{z})-d(D^{\prime},\mathbf{z})\right|$
$=$ $\left|\left\lVert\frac{1}{s}\sum_{\mathbf{z}^{\prime}\in Z}p_{\mathbf{z}^{\prime}}-p_{\mathbf{z}}\right\rVert_{1}-\left\lVert\frac{1}{s}\sum_{\mathbf{z}^{\prime}\in Z^{\prime}}p_{\mathbf{z}^{\prime}}-p_{\mathbf{z}}\right\rVert_{1}\right|$
$\leq$ $\left\lVert\frac{1}{s}\sum_{\mathbf{z}^{\prime}\in Z}p_{\mathbf{z}^{\prime}}-\frac{1}{s}\sum_{\mathbf{z}^{\prime}\in Z^{\prime}}p_{\mathbf{z}^{\prime}}\right\rVert_{1}$
$=$ $\left\lVert\frac{1}{s}\mathbf{p}_{\mathbf{\tilde{z}}}\right\rVert_{1}$
$=$ $\frac{1}{s}$
where we used the reverse triangle inequality. ∎
C.4 Constituent mechanisms
In this section we prove privacy guarantees for several simpler mechanisms that we will later compose together to show that Algorithm 1 is private.
Both Algorithms 2 and 3 accept a sensitive prompt dataset and a token sequence as input. Algorithm 2 appends a private token to the token sequence, while Algorithm 3 appends zero or more public tokens to the token sequence. The operation of both algorithms is governed by the parameters of Algorithm 1 (e.g., temperature, noise level, etc).
Algorithm 2 Private token generation
1:Input: Sensitive prompt dataset $D$, initial token sequence $\mathbf{x}_{0}$
2:Output: Token sequence $\mathbf{x}\in\mathcal{X}^{*}$
3:$\mathbf{x}\leftarrow\mathbf{x}_{0}$
4:$Z\leftarrow\{\operatorname{logits}(\mathbf{p}\mathbf{x}):\mathbf{p}\in D\}$
5:$\bar{\mathbf{z}}\leftarrow\ell(Z)$
6:$x\sim\operatorname{softmax}(\bar{\mathbf{z}}/\tau)$
7:Append $x$ to $\mathbf{x}$
8:Return $\mathbf{x}$.
Lemma 4.
Let $A(D,\mathbf{x}_{0})$ be Algorithm 2. For each $\mathbf{x}_{0}\in\mathcal{X}^{*}$ the mechanism $M:D\mapsto A(D,\mathbf{x}_{0})$ satisfies $\rho$-zCDP, where $\rho=\frac{1}{2}(\frac{c}{s\tau})^{2}$.
Proof.
Consider a function $f:\mathcal{D}\rightarrow\mathbb{R}^{v}$ with sensitivity $\Delta$. By an analysis of the exponential mechanism due to Cesar and Rogers [2021], [^6] See also Rogers and Steinke [2021]. choosing a token according to the distribution $\operatorname{softmax}(\frac{\varepsilon}{2\Delta})$ satisfies $\frac{1}{8}\varepsilon^{2}$-zCDP. Observe that mechanism $M$ is the exponential mechanism with $f=\frac{1}{\tau}\ell$, which by Lemma 3 has sensitivity $\frac{c}{s\tau}$. ∎
Algorithm 3 Public token generation
1:Input: Sensitive prompt dataset $D$, initial token sequence $\mathbf{x}_{0}$
2:Output: Token sequence $\mathbf{x}\in\mathcal{X}^{*}$
3:$\mathbf{x}\leftarrow\mathbf{x}_{0}$
4:$\hat{\theta}\leftarrow\theta+\textrm{Laplace}(\sigma)$
5:while True do
6:     $Z\leftarrow\{\operatorname{logits}(\mathbf{p}\mathbf{x}):\mathbf{p}\in D\}$
7:     $\mathbf{z}_{\operatorname{public}}\leftarrow\operatorname{logits}(\mathbf{p}_{\operatorname{public}}\mathbf{x})$
8:     $\hat{d}\leftarrow d(Z,\mathbf{z}_{\operatorname{public}})+\textrm{Laplace}(2\sigma)$
9:     if $\hat{q}>\hat{\theta}$ then
10:         Break
11:     else
12:         $x\sim\operatorname{softmax}(\mathbf{z}_{\operatorname{public}}/\tau_{\operatorname{public}})$
13:         Append $x$ to $\mathbf{x}$
14:Return $\mathbf{x}$.
Lemma 5.
Let $A(D,\mathbf{x}_{0})$ be Algorithm 3. For each $\mathbf{x}_{0}\in\mathcal{X}^{*}$ the mechanism $M:D\mapsto A(D,\mathbf{x}_{0})$ satisfies $\rho$-zCDP, where $\rho=\frac{2}{(s\sigma)^{2}}$.
Proof.
Observe that mechanism $M$ is an instance of the AboveThrehold mechanism [Dwork et al., 2009], which accepts a private dataset, a threshold, and a sequence of queries as input. In each iteration, the AboveThreshold mechanism applies the next query in the sequence to the dataset and compares it to a noisy threshold, and returns the index of the first query that exceeds the threshold. The queries can be chosen adaptively and adversarially. In mechanism $M$, each query is specified by a token sequence $\mathbf{x}$, and the index of the first query that exceeds the threshold is determined by the length of the returned token sequence. Furthermore, by Lemma 3 each query has sensitivity $\frac{1}{s}$. Thus by the analysis due to Dwork et al. [2009], mechanism $M$ satisfies $\frac{2}{s\sigma}$-differential privacy, which by Bun and Steinke [2016] implies that mechanism $M$ satisfies $\frac{2}{(s\sigma)^{2}}$-zCDP. ∎
C.5 Putting it all together
Consider a sequence of iterations of the inner loop of Algorithm 1 in which the value of $t$ (the private token counter) is constant. Observe that the operation of Algorithm 1 during these iterations is equivalent to the sequential composition of Algorithms 2 and 3, since these iterations generate zero or more public tokens followed by a private token. [^7] The special treatment of the <eos> token complicates this story a little, but we can always assume that the LLM ignores any tokens before the last <eos> token. Moreover, there are at most $r$ such sequences of iterations, since $r$ is an upper bound on the private token counter for any batch. By Lemmas 1, 4 and 5 we have that Algorithm 1 applied to a single batch satisfies $\rho$-zCDP (where $\rho$ is specified in the statement of Theorem 1). And therefore by Assumption 1 and Lemma 2 we have that Algorithm 1 applied to the whole dataset satisfies $\rho$-zCDP. It remains to convert this zCDP guarantee to an $(\varepsilon,\delta)$-differential privacy guarantee, which we do two different ways using two existing results: Corollary 13 due to Canonne et al. [2020] and Lemma 3.5 due to Bun and Steinke [2016].
Appendix D Privacy-equivalent Gaussian noise
Given the average token distribution $\bar{\mathbf{p}}$ in a batch, Tang et al. [2024] protect the privacy of $\bar{\mathbf{p}}$ by using the Gaussian mechanism, which achieves $(\varepsilon,\delta)$-differential privacy with $\varepsilon=\frac{\sqrt{2\log(1.25/\delta)}}{s\sigma}$, where $s$ is the batch size and $\sigma$ is the standard deviation of the noise added to each probability in $\bar{\mathbf{p}}$. On the other hand, we use the exponential mechanism to protect the privacy of a sample drawn from $\bar{\mathbf{p}}$, which achieves $\varepsilon$-differential privacy with $\varepsilon=\frac{2c}{s\tau}$, where $c$ is the maximum absolute value of any log-probability in the batch and $\tau$ is the sampling temperature.
Empirically, we obtained good synthetic data quality with $s=250$, $\tau=2$, $c=10$ and $\delta=10^{-6}$.
Setting the $\varepsilon$ values equal to each other yields $\sigma=\frac{\tau\sqrt{\log(1.25/\delta)}}{\sqrt{2}c}$, which is the noise level needed for the two mechanisms to have comparable privacy guarantees (setting aside that $\delta>0$, an omission that only favors the Gaussian mechanism). Plugging in the above parameters yields $\sigma\approx 0.53$.
The analysis in Theorem 8 of Balle and Wang [2018] does not admit a closed-form solution. Instead, we binary search for a solution to:
$\Phi\left(\frac{\Delta}{2\sigma}-\frac{\varepsilon\sigma}{\Delta}\right)-\exp(\varepsilon)\Phi\left(-\frac{\Delta}{2\sigma}-\frac{\varepsilon\sigma}{\Delta}\right)\leq\delta$
where $\Phi$ is the Gaussian cdf, $\varepsilon=\frac{2c}{s\tau}$, $\delta=10^{-6}$, and $\Delta$ is the L2 sensitivity of a vector computed as the average of $s$ user-provided probability vectors, namely $\Delta=1/s$. This procedure yields $\sigma\approx 0.34$.
Finally, equating the zCDP loss for the exponential mechanism given by $\frac{\varepsilon^{2}}{8}=\frac{c^{2}}{2s^{2}\tau^{2}}$ (Cesar and Rogers [2021]) to that of the Gaussian mechanism given by $\frac{1}{2s^{2}\sigma^{2}}$ (Bun and Steinke [2016]), yields $\sigma=0.2$.
Appendix E Experiment details
E.1 Hyperparameter tuning
There are a significant amount of hyperparameters associated with our approach. See Table 6 for a list of the main ones and the values they take. In this section we describe the hyperparameter evaluation procedure, as well as the rationale for our decisions on what hyperparameter settings to couple together or that we altogether avoid running.
Hyperparameter evaluation procedure.
For fine-tuning experiments, we set aside a real validation set consisting of 10% the real train set. We choose dataset generation parameters based on which resulting dataset induces the the best classifier on this real validation set. However, the process of tuning the classifier itself on synthetic data (choosing the best learning rate and checkpoint) does not use real data – we do that tuning with synthetic data. This is because the output of our method is a dataset, and its usefulness to train a model includes how well subsets of it can be used for downstream task hyperparameter selection. After identifying the best synthetic dataset in this manner, we run the tuning process based on synthetic data only and report accuracy of the resultant classifier on the real test set.
Hyperparameter choices.
Based on initial experiments, we found that setting $c=10$ and $\tau=2$ produced well formed text, so we fix $c=10$ and try a low temperature ($\tau=1.5$) and a high temperature ($\tau=2.25$) setting. At $\tau=2.25$, we observed text degeneration. This is due to the combination of Gemma’s large vocabulary (256K) and clipping, which raises the “probability floor” of nonsense tokens. So for $\tau=2.25$ settings only, we follow Tang et al. [2024] and reduce the vocabulary to the public prediction’s top 1024. We emphasize that (1) we do not do this for any of the other settings of $\tau$, and (2) use a larger value than prior work (they use top 100).
Keeping other parameters fixed and increasing the batch size $s$ decreases $\varepsilon$. At the same time, it raises the amount of compute spent to decode a single example. [^8] The way we interpret this is that $s$ is a compute multiplier that broadens the search space to include better utility configurations in the low $\varepsilon$ regime. This is analagous to the role of the noise multiplier $\sigma$ in DP-SGD, where the best results at low $\varepsilon$ come from taking more steps at higher noise levels. Hence our approach for selecting the batch size is based on the following: given a target epsilon and dataset, choose $s$ large enough so that we can hit at least 1K examples at the low temperature setting $\tau=1.5$. When targeting a large $\varepsilon$, choosing large $s$ results in too many tokens to decode at too high of a cost per token.
[…]
$\alpha$ Description Values
$s$ batch size 127, 255, 511,
1023, 1535, 2047
$c$ logits clip bound 10
$\tau$ temperature 1.5, 2, 2.25
$\theta$ SVT threshold $\infty$- ∞, 0.3, 0.5, 0.7
$\sigma$ SVT noise level --, 0.1, 0.2, 0.2
$\tau_{\text{public}}$ public temperature 1.5
Table 6: Values for hyperparameters explored in this work.
E.2 Prompts used
We report the prompts used for our experiments. Generally, we use the same prompt for private and public predictions, with "<text of xxx>" in the public prompt replaced with an actual private example in the private prompt. The exception is for WikiMoviesJSON (Figures 11 and 12), where the public prompt contains a schema description in place of the example.
1# [User] 2Here are texts with News Type: Business. 3 4Text: <text of News Type: Business> 5 6Please give me another one. 7 8# [Assistant] 9Text:
Figure 4: Generation prompt for AGNews.
1# [User] 2Here are questions with Answer Type: Entity. 3 4‘‘‘ 5Text: <question of Answer Type: Entity> 6‘‘‘ 7 8Please give me another one. 9 10# [Assistant] 11‘‘‘ 12Question:
Figure 5: Generation prompt for TREC.
1# [User] 2Here are entries of Category: School. 3 4Entry: <entry of Category: School> 5 6Please give me another one. 7 8# [Assistant] 9Entry:
Figure 6: Generation prompt for DBPedia.
1# [User] 2Give me text about a film and the extracted Phrase about its Director. 3 4Phrase: "josh trank" 5Text: "<text containing phrase "josh trank">" 6 7Please give me another Phrase and Text: "josh trank". IMPORTANT: The exact Director phrase "josh trank" must be mentioned in Text. 8 9# [Assistant] 10Phrase: "josh trank" 11Text: ""
Figure 7: Generation prompt for MIT-D.
1# [User] 2Give me text about a film and the extracted Phrase about its Genre. 3 4Phrase "comedy" 5Text: "<text containing phrase "comedy">" 6 7Please give me another Phrase and Text. IMPORTANT: The exact Genre phrase "comedy" must be mentioned in Text. 8 9# [Assistant] 10Phrase: "comedy" 11Text: ""
Figure 8: Generation prompt for MIT-G.
1# [User] 2Here are texts with Sentiment: Negative. 3 4Text: <text of Sentiment: Negative> 5 6Please give me another one. 7 8# [Assistant] 9Text:
Figure 9: Generation prompt for IMDB.
1# [User] 2Here are texts with Sentiment: Negative. 3 4Text: <text of Sentiment: Negative> 5 6Please give me another one. 7 8# [Assistant] 9Text:
Figure 10: Generation prompt for Yelp.
1# [User] 2Here is a JSON record: 3‘‘‘ 4{ 5 "title": "\$50,000 Reward", 6 "year": 1924, 7 "cast": [ 8 "Ken Maynard", 9 "Esther Ralston" 10 ], 11 "genres": [ 12 "Western", 13 "Silent" 14 ], 15 "href": "\$50,000_Reward", 16 "extract": "\$50,000 Reward is a 1924 American silent Western film directed by Clifford S. Elfelt and starring Ken Maynard, Esther Ralston and Bert Lindley." 17} 18‘‘‘ 19Please give me another JSON record that complies with the above schema. 20 21# [Assistant] 22‘‘‘ 23{
Figure 11: Private generation prompt for WikiMoviesJSON.
1# [User] 2Here is the schema for a JSON record: 3Schema: 4‘‘‘ 5{ 6 "title": "str", 7 "year": int, 8 "cast": [ # list of str 9 "str1", # 0 or more total entries 10 ], 11 "genres": [ # list of str 12 "str1", # 0 or more total entries 13 ] 14 "href": "str", # URL slug, e.g.: Link_to_Page 15 "extract": "str" 16} 17‘‘‘ 18Please give me another JSON record that complies with the above schema. 19 20# [Assistant] 21‘‘‘ 22{
Figure 12: Public generation prompt for WikiMoviesJSON.
Appendix F Artifacts
Tables 1(a) and 1(b) list all artifacts we use in this work. AGNews, TREC, DBPedia, MIT-G, MIT-D, IMDB, and Yelp are all standard academic datasets permissible for research use; we cite their original publications when introduced. WikiMoviesJSON is scraped from Wikipedia data, courtesy of [Rust, 2024]; their work is covered by an MIT license. Wikipedia content is licensed under the Creative Commons Attribution-ShareAlike 4.0 International License (CC BY-SA) and the GNU Free Documentation License (GFDL).
We use open-source models BERT-Base, released by [Turc et al., 2019], and Gemma. Our use of Gemma for academic purposes is in accordance of the Gemma terms of use: https://ai.google.dev/gemma/terms. GPT-3 is accessible for academic purposes under OpenAI’s terms of use, which supports educational and research activities. LaMDA 8B is not publically available, but we received sufficient authorization to use it for the academic purposes of this paper.
Appendix G Compute budget
All experiments for synthetic data generation run on Gemma 2B 1.1 IT. A run of synthetic data generation takes between 8-48 accelerator hours. Including exploratory runs and hyperparameter search, the total compute budget for this project is roughly 14,000 accelerator hours.

### v2

@@
-Kareem Amin  Alex Bie  Weiwei Kong  Alexey Kurakin Natalia Ponomareva  Umar Syed  Andreas Terzis  Sergei Vassilvitskii Google
+Private prediction for large-scale synthetic text generation
 Abstract
@@
 Furthermore, we introduce a novel use of public predictions via the sparse vector technique, in which we do not pay privacy costs for tokens that are predictable without sensitive data; we find this to be particularly effective for structured data.
+Private prediction for large-scale synthetic text generationthanks: Authors ordered alphabetically. Author contributions are listed at the end.
+Kareem Amin Alex Bie Weiwei Kong Alexey Kurakin Natalia Ponomareva Umar Syed Andreas Terzis Sergei Vassilvitskii Google {kamin,alexbie,weiweikong,kurakin,nponomareva,usyed,aterzis,sergeiv}@google.com
 1 Introduction
@@
 However, unlike private fine-tuning, the guarantees of private prediction methods degrade with the volume of data that is generated.
-Consequently, existing private prediction methods have mostly been used in applications that require only small amounts of synthetic data [Tang et al., 2024], sharply limiting their practicality.
+Consequently, existing private prediction methods have mostly been used in applications that require only small amounts of synthetic data (Tang et al., 2024), sharply limiting their practicality.
+Figure 1: Algorithm 1, visualized.
+An LLM receives a batch of prompts, each instructing to generate text similar to a piece of sensitive text.
+Synthetic text is generated token by token, by running inference on the batch in parallel.
+In each step, the logit vectors produced downstream of sensitive text are aggregated and sampled from with differential privacy.
+Every token sampled in such way incurs a privacy cost, motivating us to include an auxillary public prompt and sample from its logits when similar to the sensitive logits.
+1.1 Our contributions
 In this paper we describe a new private prediction method that produces hundreds of times as much synthetic data as a state-of-the-art private prediction method, while maintaining a comparable privacy guarantee.
@@
 However, our approach is distinguished by three novel algorithmic elements that lead to its improved performance:
-  1. 1.
+1. Better private token selection.
 Instead of protecting the privacy of the entire token distribution with Gaussian or Laplace noise, we leverage the uncertainty inherent in sampling to ensure privacy.
-We clip and aggregate token logits before standard softmax sampling — which is differentially private, since it can viewed as the exponential mechanism.
+We clip and aggregate token logits before standard softmax sampling — which is differentially private, since it can be viewed as the exponential mechanism.
 Our approach induces much less distortion of the original token distributions to achieve the same level of privacy than prior work.
-  2. 2.
-    Previous work generated each token using a random subset of the input data, leveraging privacy amplification by subsampling in their analysis.
-This is computationally undesirable, as it requires repeated re-computation of the prefix for each decoding step, and limits scalability towards generating large synthetic corpora.
-Instead, we generate each synthetic example using a fixed disjoint subset of the input data, which yields substantial savings in privacy cost (by leveraging parallel composition) while allowing us to pay a linear amount of non-attention FLOPs, rather than quadratic, in terms of sequence length (via KV cache accelerated decoding).
-  3. 3.
+2. Avoiding prefix re-sampling.
+Prior work generated each token using a random subset of the input data, leveraging privacy amplification by subsampling in their analysis.
+This is computationally undesirable, as it requires repeated re-computation of the prefix for each decoding step, thus limiting scalability towards generating large synthetic corpora.
+Indeed, prior work describes this re-sampling as the “main weakness” of the approach Tang et al.
+(2024).
+To resolve the problem, we instead generate each synthetic example using a fixed disjoint subset of the input data, which yields substantial savings in privacy cost – via parallel composition – while allowing us to pay linear instead of quadratic non-attention FLOPs in terms of sequence length via KV cache accelerated decoding.
+3. Leveraging public predictions.
 Our method uses an auxillary token distribution from an LLM without access to sensitive data, and draws the next token from that distribution whenever it is very similar to the token distribution induced by the sensitive data.
@@
 Taken together, the combination of these algorithmic techniques leads to significant improvements over prior work.
-Informally, (1) and (2) above keep our inference closely aligned to standard (non-DP) inference.
+Roughly speaking, (1) and (2) above keep our inference closely aligned to standard (non-DP) inference.
 In our experiments, we generate private synthetic versions of publicly available, benchmark machine learning datasets, and then use the synthetic datasets for downstream classification and extraction tasks.
 Owing to the increased quantity and quality of our synthetic data, we improve over an existing state-of-the-art private prediction method in terms of downstream accuracy.
-Furthermore, while prior work in this paradigm only generated a small (<10) number of examples we demonstrate the ability to generate thousands of training examples, enough for fine-tuning downstream models.
-Finally, since synthetic data is intended for a wide variety of applications, we also evaluate data quality using a metric that is orthogonal to performance on a downstream classification task.
+Furthermore, while prior work in this paradigm only generated a small (<10) number of examples, we demonstrate the ability to generate thousands of training examples, enough for fine-tuning downstream models.
+Finally, since synthetic data is intended for a wide variety of applications, we also evaluate data quality using a metric that is orthogonal to performance on downstream tasks.
 Specifically, we generate synthetic versions of a publicly available dataset containing highly structured data records, each of which is encoded as a JSON object.
@@
 Yue et al.
-[2023] created private synthetic versions of text datasets by using DP-SGD [Abadi et al., 2016] to fine-tune an LLM on the sensitive data.
+(2023) created private synthetic versions of text datasets by using DP-SGD (Abadi et al., 2016) to fine-tune an LLM on the sensitive data.
 Kurakin et al.
-[2024] showed that parameter efficient approaches to fine-tuning, such as LoRA [Hu et al., 2022], can improve the quality of the synthetic data, since reducing the number of parameters also reduces the amount of noise injected into the optimization procedure.
+(2024) showed that parameter efficient approaches to fine-tuning, such as LoRA (Hu et al., 2022) can improve the quality of the synthetic data.
 Wu et al.
-[2024a] took a two-stage approach: First they fine-tuned an LLM on a public dataset that closely resembled the sensitive data (which was itself generated by an LLM using carefully designed prompts), and then completed the fine-tuning process by running DP-SGD on the sensitive dataset.
-Concurrent to this work, Tran and Xiong [2024] describe a private fine-tuning approach for generating synthetic tabular data that is formatting compliant.
-Private prediction [Dwork and Feldman, 2018] is an alternate approach to private machine learning that only guarantees the privacy of the predictions output by an ML model, and not the model itself.
+(2024a) took a two-stage approach: First they fine-tuned an LLM on a public dataset that closely resembled the sensitive data (which was itself generated by an LLM using carefully designed prompts); then they completed the fine-tuning process by running DP-SGD on the sensitive dataset.
+Concurrent to the present work, Tran and Xiong (2024) describe a private fine-tuning approach for generating synthetic tabular data that is formatting compliant.
+Private prediction (Dwork and Feldman, 2018) is an alternate approach to private machine learning that only guarantees the privacy of the predictions output by an ML model, and not the model itself.
+The predominant way this is realized is via subsample-and-aggregate (Nissim et al., 2007): First sensitive data is split into disjoint partitions; then non-private predictions are made from each partition and privately aggregated.
+PATE (Papernot et al., 2017, 2018) employs this approach to get answers to a limited set of image classification queries, which are then used to train a student model that can be queried indefinitely.
 Private prediction has been applied to synthetic text generation by viewing each token sampled by an LLM as a ‘prediction’, and perturbing the LLM’s token distributions to ensure their privacy.
 Tang et al.
-[2024] added noise to several independent token distributions and averaged them, while Hong et al.
-[2024] selected the most popular token among the token distributions using the LimitedDomain mechanism [Durfee and Rogers, 2019].
+(2024) added noise to several independent token distributions and averaged them, while Hong et al.
+(2024) selected the most popular token among the token distributions using the LimitedDomain mechanism (Durfee and Rogers, 2019).
 These methods can avoid the time, compute, and access required to fine-tune an LLM with billions of parameters.
@@
 As a result, previous work has only been able to generate a very small number of synthetic examples at reasonable privacy levels (fewer than 10).
-Other work has applied private prediction techniques to LLMs [Majmudar et al., 2022, Duan et al., 2023], including in combination with fine-tuning [Ginart et al., 2022, Flemings et al., 2024], but not for the purpose of synthetic text generation.
+Other work has applied private prediction techniques to LLMs (Majmudar et al., 2022; Duan et al., 2023), including in combination with fine-tuning (Ginart et al., 2022; Flemings et al., 2024), but not for the purpose of synthetic text generation.
 Finally, another distinct set of approaches are private filtering methods.
@@
 Yu et al.
-[2024] and Xie et al.
-[2024] used the sensitive responses to privately select similar responses from the public dataset.
+(2024) and Xie et al.
+(2024) used the sensitive responses to privately select similar responses from the public dataset.
 Similarly, Wu et al.
-[2024b] aggregate response embeddings and select the public response that is closest in embedding space.
+(2024b) aggregate response embeddings and select the public response that is closest in embedding space.
 [^1] Wu et al.
-[2024b] also proposed a non-filtering approach based on privately selecting common keywords among the sensitive data and using them to prompt an LLM.
+(2024b) also proposes a non-filtering approach based on privately selecting common keywords among the sensitive data and using them to prompt an LLM.
 One limitation of filtering methods is that the menu of possible responses is constructed without signal from the new source dataset.
 3 Method
+3.1 Standard LLM inference
 Before describing our algorithm for generating private synthetic text, we review the standard algorithm for LLM inference.
@@
 If $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$ are token sequences then we write $\mathbf{x}_{1}\mathbf{x}_{2}\in\mathcal{X}^{*}$ to denote their concatenation.
-Standard LLM inference.
 […]
-Our algorithm.
+3.2 Our algorithm
 One straightforward approach to generating a synthetic version of a sensitive piece of text would be to prompt an LLM with ‘Please generate text similar to: <sensitive text>’.
 However, this could easily lead to a privacy violation, as the response could retain the semantics of the input sensitive text.
-Figure 1: Visualization of Algorithm 1 for a single token in a single batch.
 Algorithm 1 describes our method for privately generating a dataset of synthetic examples $X$ from a dataset of sensitive prompts $D$.
 Each prompt in $D$ resembles the sample prompt given above.
-But instead of using a single prompt to generate a synthetic example, the algorithm uses a batch of the prompts to run several LLM inferences in parallel.
-Each synthetic example is generated one token at a time, with the average of the logit vectors across the inferences defining the distribution from which the next token is randomly selected.
-Before averaging, the logits are clipped and re-centered using the function
-$\operatorname{clip}_{c}(\mathbf{z})_{i}=\max\left(-c,\mathbf{z}_{i}-\max_{j}{\{\mathbf{z}_{j}\}}+c\right)$ (1)
-which maps each component of $\mathbf{z}$ into $[-c,c]$.
+But instead of using a single prompt to generate a synthetic example, Algorithm 1 takes a batch of prompts $S$ and runs LLM inference in parallel on each prompt.
+A synthetic example is generated one token at a time, with the average of the logit vectors across the batch defining the distribution from which the next token is randomly selected.
+Before averaging, a logit vector $\mathbf{z}$ is clipped and re-centered using the function
+$\operatorname{clip}_{c}(\mathbf{z})_{i}=\max\{-c,\mathbf{z}_{i}-\max_{j}{\{\mathbf{z}_{j}\}}+c\}$ (1)
+which maps each component $i$ of $\mathbf{z}$ into the target clipping range $[-c,c]$.
 Forcing each logit to lie in a bounded range is key to proving the privacy guarantee for our algorithm (see §4).
@@
 This property is desirable because the distribution $\operatorname{softmax}(\mathbf{z})$ is invariant to any constant shift of $\mathbf{z}$.
-We also found that Eq.
-(1) performed better empirically than other functions we considered.
-For example, regular clipping to the range $[-c,c]$ without recentering requires twice as large $c$ to sample without distortion (see Appendix B).
-Algorithm 1 Generate private synthetic examples using an LLM
-1:Parameters: LLM logit function $\operatorname{logits}(\cdot)$, public prompt $\mathbf{p}_{\operatorname{public}}$, expected batch size $s$, maximum number of private tokens per batch $r$, clipping value $c$, noise level $\sigma$, distance function $d$, threshold $\theta$, public temperature $\tau_{\operatorname{public}}$, private temperature $\tau$
-2:Input: Dataset of sensitive prompts $D$
+Empirically, we found that Eq.
+(1) performed better than other functions considered.
+For example, regular clipping to the range $[-c,c]$ without recentering requires twice as large $c$ to sample without distortion (see §B.1).
+Algorithm 1 Generate private synthetic examples
+1:Parameters: LLM $\operatorname{logits}(\cdot)$, public prompt $\mathbf{p}_{\operatorname{public}}$, expected batch size $s$, private tokens to sample $r$.
+Sampling: clipping bound $c$, temperature $\tau$, public temperature $\tau_{\operatorname{public}}$.
+Public optionality: public/private distribution distance $\mathrm{d}(\cdot,\cdot)$, threshold $\theta$, noise level $\sigma$.
+2:Input: Dataset of sensitive prompts $D$; each prompt contains a sensitive example
 3:Output: Dataset of synthetic examples $X$
-4:Let $\mathcal{S}$ be a partition of $D$ into disjoint batches
-5:for each batch $S\in\mathcal{S}$ do
-6:     $\hat{\theta}\leftarrow\theta+\textrm{Laplace}(\sigma)$
-7:     $t\leftarrow 0$ private token counter
-8:     while $t<r$ do
+4:$X\leftarrow\emptyset$
+5:Let $\mathcal{S}$ be a partition of $D$ into disjoint batches
+6:for each batch $S\in\mathcal{S}$ do
+7:     $\hat{\theta}\leftarrow\theta+\textrm{Laplace}(\sigma)$# Init noisy threshold
+8:     $t\leftarrow 0$# Private token counter
+9:     while $t<r$ do
 […]
-10:         while $\mathbf{x}$ does not end with <eos> do
-11:              $Z\leftarrow\{\operatorname{logits}(\mathbf{p}\mathbf{x}):\mathbf{p}\in S\}$
-12:              $\mathbf{z}_{\operatorname{public}}\leftarrow\operatorname{logits}(\mathbf{p}_{\operatorname{public}}\mathbf{x})$
-13:              $\hat{d}\leftarrow d(Z,\mathbf{z}_{\operatorname{public}})+\textrm{Laplace}(2\sigma)$
-14:              if $\hat{d}>\hat{\theta}$ then
-15:                  $\bar{\mathbf{z}}\leftarrow\frac{1}{s}\sum_{\mathbf{z}\in Z}\operatorname{clip}_{c}(\mathbf{z})$
-16:                  $x\sim\operatorname{softmax}(\bar{\mathbf{z}}/\tau)$
-17:                  $\hat{\theta}\leftarrow\theta+\textrm{Laplace}(\sigma)$
+11:         while $\mathbf{x}$ does not end with <eos> do
+12:              $Z\leftarrow\{\operatorname{logits}(\mathbf{p}\mathbf{x}):\mathbf{p}\in S\}$
+13:              $\mathbf{z}_{\operatorname{public}}\leftarrow\operatorname{logits}(\mathbf{p}_{\operatorname{public}}\mathbf{x})$     # Check if pub/priv distributions are far
+14:              $\hat{d}\leftarrow\mathrm{d}(Z,\mathbf{z}_{\operatorname{public}})+\textrm{Laplace}(2\sigma)$
+15:              if $\hat{d}\geq\hat{\theta}$ then# Sample priv token
+16:                  $\bar{\mathbf{z}}\leftarrow\frac{1}{s}\sum_{\mathbf{z}\in Z}\operatorname{clip}_{c}(\mathbf{z})$
+17:                  $x\sim\operatorname{softmax}(\bar{\mathbf{z}}/\tau)$
 18:                  $t\leftarrow t+1$
-19:              else
-20:                  $x\sim\operatorname{softmax}(\mathbf{z}_{\operatorname{public}}/\tau_{\operatorname{public}})$
-21:              Append $x$ to $\mathbf{x}$
-22:         Add $\mathbf{x}$ to $X$
-23:Return $X$.
-Since the average logit vector is computed using a set of sensitive prompts, each token selected from a distribution determined by the average logit vector incurs a privacy cost.
+19:                  $\hat{\theta}\leftarrow\theta+\textrm{Laplace}(\sigma)$
+20:              else# Sample pub token
+21:                  $x\sim\operatorname{softmax}(\mathbf{z}_{\operatorname{public}}/\tau_{\operatorname{public}})$
+22:              Append $x$ to $\mathbf{x}$
+23:         $X\leftarrow X\cup\{\mathbf{x}\}$
+24:return $X$
+Since the average logit vector is computed using sensitive prompts, each token selected from a distribution determined by the average logit vector incurs a privacy cost.
 To minimize this cost, Algorithm 1 also has access to a non-sensitive public prompt, $\mathbf{p}_{\operatorname{public}}$, and uses this prompt to generate the next token whenever doing so does not significantly change the distribution from which the next token is drawn.
 The distance function used to make this determination is
-$d(Z,\mathbf{z}_{\operatorname{public}})=\left\lVert\frac{1}{s}\sum_{\mathbf{z}\in Z}p_{\mathbf{z}}-p_{\mathbf{z}_{\operatorname{public}}}\right\rVert_{1}$ (2)
-where $p_{\mathbf{z}}=\operatorname{softmax}(\mathbf{z})$ is the token distribution corresponding to logit vector $\mathbf{z}$.
+$\mathrm{d}(Z,\mathbf{z}_{\operatorname{public}})=\|\frac{1}{s}\sum_{\mathbf{z}\in Z}p_{\mathbf{z}}-p_{\mathbf{z}_{\operatorname{public}}}\|_{1},$ (2)
+where $p_{\mathbf{z}}:=\operatorname{softmax}(\mathbf{z})$, $Z$ are the logit vectors computed for each sensitive prompt in $S$, $\mathbf{z}_{\operatorname{public}}$ is the logit vector computed using $\mathbf{p}_{\operatorname{public}}$, and $s$ is the expected batch size.
 When this distance is small, Algorithm 1 outputs a public token instead of a private token.
-The privacy guarantee for Algorithm 1 leverages the analysis of the sparse vector technique [Dwork et al., 2009], and shows that while privacy degrades with the number of private output tokens, it is independent of the number of public output tokens (see §4).
-Empirically, we observe that the fraction of output tokens that must be private in order to generate high-quality synthetic data can be only 20% for some datasets.
+The privacy guarantee for Algorithm 1 leverages the analysis of the sparse vector technique (Dwork et al., 2009), and shows that while privacy degrades with the number of private output tokens, it is independent of the number of public output tokens (see §4).
+Empirically, we observe that the fraction of output tokens that must be private in order to generate high-quality synthetic data can be as low as 20% for highly structured datasets.
 Note that the first step of Algorithm 1 partitions the input dataset of sensitive prompts into disjoint batches.
 We do not prescribe a procedure for assigning prompts to batches in Algorithm 1 since many batching approaches are admissible as long as they satisfy a minor technical assumption required for the privacy analysis of Algorithm 1, which we explain in §4.
-While the batches are not required to be any particular size, the algorithm makes most efficient use of the data if each batch has size equal to the expected batch size $s$.
+While the batches are not required to be any particular size, the algorithm runs faster if each batch has size equal to the expected batch size $s$.
 And while prompts can be batched together (almost) arbitrarily, more tailored batching can lead to better synthetic data quality.
@@
 In those experiments we assign prompts with the same label to the same batch.
-Relationship to prior work.
+3.3 Comparison to prior algorithms
 Two major features of Algorithm 1 are that it leverages the inherent randomness of token sampling to guarantee privacy, and that it further reduces privacy cost by using public data to generate a portion of the synthetic data.
@@
 Instead of clipping logits to ensure that the token sampling is private, Majmudar et al.
-[2022] mixed each sensitive token distribution with the uniform distribution.
+(2022) mixed each sensitive token distribution with the uniform distribution.
 This approach induced a dependence on the vocabulary size in their privacy guarantee, and since LLM vocabularies are typically very large, the resulting privacy guarantee was quite weak: Majmudar et al.
-[2022] noted that setting the differential privacy parameter $\varepsilon$ (see Definition 1) lower than 50 produced synthetic data that was “unusable”.
+(2022) noted that setting the differential privacy parameter $\varepsilon$ lower than 50 produced synthetic data that was “unusable”.
 Flemings et al.
-[2024] guaranteed the privacy of token sampling by mixing each sensitive token distribution with a public token distribution, but their approach was based on aggregating a set of fine-tuned models, not a set of prompts.
+(2024) guaranteed the privacy of token sampling by mixing each sensitive token distribution with a public token distribution, but their approach was based on aggregating a set of fine-tuned models, not a set of prompts.
 Neither Majmudar et al.
-[2022] nor Flemings et al.
-[2024] had a goal of generating synthetic data.
+(2022) nor Flemings et al.
+(2024) aim to generate synthetic data.
 Tang et al.
-[2024] found that limiting the token vocabulary to a fixed set of the most popular 100 public tokens caused their synthetic data generation algorithm to exhibit greater stability.
+(2024) found that limiting the token vocabulary to a fixed set of the most popular 100 public tokens caused their synthetic data generation algorithm to exhibit greater stability.
 However, if the sensitive data contains many tokens that are rare in public data, their approach cannot produce synthetic data that is very similar to the sensitive data.
@@
 Also, Tang et al.
-[2024] used a different random subset of prompts to generate each token, and left as an open problem how to use a single subset to generate every token in a synthetic example.
+(2024) used a different random subset of prompts to generate each token, and left as an open problem how to use a single subset to generate every token in a synthetic example.
 Our algorithm resolves this open problem, and consequently yields both improved privacy and greater computational efficiency (see §6).
 4 Privacy analysis
-In this section we prove that Algorithm 1 preserves the privacy of the sensitive prompts it uses to generate synthetic examples.
+In this section we state formally how Algorithm 1 preserves the privacy of the sensitive prompts it uses to generate synthetic examples.
 Let $\mathcal{D}$ be the set of all possible prompt datasets.
@@
 In the differential privacy literature this is commonly referred to as the add/remove neighbor relation.
-Definition 1 (Dwork et al. [2006]).
+Definition 1 (Dwork et al. (2006)).
 A mechanism $M$ satisfies $(\varepsilon,\delta)$-differential privacy if $\Pr[M(D)\in O]\leq e^{\varepsilon}\Pr[M(D^{\prime})\in O]+\delta$ for any neighboring datasets $D,D^{\prime}\in\mathcal{D}$ and subset $O$ of the range of $M$.
@@
 $\varepsilon=\rho+\sqrt{4\rho\log(1/\delta)}.$
-The proof is in Appendix C and makes use of sharp privacy analyses of: (1) zCDP to approximate DP conversion [Canonne et al., 2020]; and (2) zCDP bounds for the exponential mechanism [Cesar and Rogers, 2021].
+The proof is in §D and makes use of sharp privacy analyses of: (1) zCDP to approximate DP conversion (Canonne et al., 2020); and (2) zCDP bounds for the exponential mechanism (Cesar and Rogers, 2021).
 5 Experiments
-Gemma 1.1 2B IT [Team, 2024] is the data generator in all private prediction experiments.
-We choose it due to its lightweight, open-source JAX implementation that makes easy to implement and share sampling algorithms.
-[^2] https://github.com/google-deepmind/gemma Tables 1(a) and 1(b) give an overview of datasets and models used.
+Gemma 1.1 2B IT (Gemma Team, 2024) is the data generator in our main private prediction experiments.
+We choose it due to its lightweight, open-source JAX implementation [^2] https://github.com/google-deepmind/gemma that makes easy to implement and share sampling algorithms.
+Tables 1(a) and 1(b) give an overview of datasets and models used.
 Dataset $n_{\text{train}}$ Description
@@
 Model Usage
-Gemma 2B 1.1 IT Generation; private prediction
+Gemma 1.1 2B IT Generation; private prediction
 LaMDA 8B Generation; DP fine-tuning
@@
 BERT-Base 12/768 110M Evaluation; fine-tuning
-(b) Overview of models used.
-Table 1: Overview of datasets and models used.
+(b) Overview of models used in main experiments.
+Table 1: Overview of datasets and models used in our main experiments.
 Datasets are benchmark classification and extraction tasks used in prior work on private synthetic text generation, with the exception of WikiMoviesJSON, which is used for structured data experiments.
@@
 We perform 3 sets of experiments, targeting various datasets and utility criteria:
-  • In-context learning (§5.1); We generate examples to use as in-context exemplars for prompting an LLM.
-We report downstream accuracy on real test examples, when prompted with synthetic data, on 3 classification tasks (AGNews [Zhang et al., 2015], DBPedia [Zhang et al., 2015], TREC [Voorhees and Tice, 2000]) and 2 extraction tasks (MIT-G, MIT-D [Liu et al., 2012]).
-  • Fine-tuning (§5.2); We generate synthetic examples to use for fine-tuning a BERT classifier.
-We report downstream accuracy on real test examples for 3 classification tasks (IMDB [Maas et al., 2011], Yelp [Zhang et al., 2015], AGNews [Zhang et al., 2015]).
-  • Structured data (§5.3); We generate examples that must adhere to structural constraints to be useful synthetic data.
-We consider a JSON generation task (WikiMoviesJSON [Rust, 2024]), evaluating structure preservation.
+In-context learning (§5.1).
+We generate examples to use as in-context exemplars for prompting an LLM.
+We report downstream accuracy on real test examples, when prompted with synthetic data, on 3 classification tasks (AGNews (Zhang et al., 2015), DBPedia (Zhang et al., 2015), TREC (Voorhees and Tice, 2000)) and 2 extraction tasks (MIT-G, MIT-D (Liu et al., 2012)).
+Fine-tuning (§5.2).
+We generate synthetic examples to use for fine-tuning a BERT classifier.
+We report downstream accuracy on real test examples for 3 classification tasks (IMDB (Maas et al., 2011), Yelp (Zhang et al., 2015), AGNews (Zhang et al., 2015)).
+Structured data (§5.3).
+We generate examples that must adhere to structural constraints to be useful synthetic data. We consider a JSON generation task (WikiMoviesJSON (Rust, 2024)), evaluating structure preservation.
 5.1 In-context learning
 Experimental setup.
-Using our method, we generate 90-1500 examples using Gemma 2B 1.1 IT.
+Using our method, we generate 90-1500 examples using Gemma 1.1 2B IT.
 We compare against real examples, and results reported in the prior work of Tang et al.
-[2024], where they generated 4-shot examples for in-context learning.
+(2024), where they generated 4-shot examples for in-context learning.
 [^3] It is no longer possible to reproduce their results, due to changes in the OpenAI API since publication: GPT-3 babbage is now deprecated, and it is no longer possible to query for top 100 logprobs, which is required by their method.
 To evaluate generated synthetic data, we put synthetic examples in the context window before querying with the real test example, as shown in Figure 2.
-1Classify the following examples: 2Text: lorem ipsum # synthetic text 1 3Answer: label 4... 5Text: sed do eiusmod # synthetic text n 6Answer: label 7 8Text: excepteur sint # test text 9Answer:
+1Classify the following examples: 2Text: lorem ipsum # synthetic text 1 3Answer: label 4# ... 5Text: sed do eiusm # synthetic text n 6Answer: label 7 8Text: excepteur si # test text 9Answer:
 Figure 2: Example of $n$-shot in-context learning evaluation for synthetic data.
-We perform this evaluation with GPT-3 babbage-002 which has a 16K context window.
-We report results on AGNews, DBPedia, TREC, MIT-G, and MIT-D using the implementation of Zhao et al.
-[2021].
-[^4] https://github.com/tonyzhaozh/few-shot-learning.
-Following the work of Tang et al.
-[2024], we enable contextual calibration [Zhao et al., 2021] for classification but not extraction tasks.
-Our evaluation setup is a best-effort reproduction of their setup, which is no longer possible to completely reproduce due to changes to OpenAI API access (see Table 2 caption for more details).
-Due to cost, we follow prior work [Bertsch et al., 2024, Ratner et al., 2023, Lu et al., 2022, Zhao et al., 2021] and opt to subsample test sets to 250 test examples.
-We run 3 seeds of label-stratified sampling of exemplars from synthetic/real data.
 GPT-3 babbage-002 Acc. (%)*
@@
 64 $84.7_{1.5}$ $92.5_{1.6}$ $50.3_{6.1}$ $56.4_{5.4}$ $89.1_{0.7}$
-Tang et al. [2024] 4 Tang et al. [2024]* GPT-3 babbage $69.3_{4.8}$ $82.3_{3.7}$ $50.6_{6.9}$ $54.4_{7.0}$
+Tang et al. (2024) 4 Tang et al. (2024)* GPT-3 babbage $69.3_{4.8}$ $82.3_{3.7}$ $50.6_{6.9}$ $54.4_{7.0}$
 Ours 4 This work Gemma 1.1 2B IT $76.8_{4.8}$ $72.3_{2.5}$ $38.8_{6.0}$ $47.7_{2.5}$ $81.7_{2.4}$
 64 $77.5_{1.8}$ $91.5_{1.7}$ $57.9_{3.4}$ $56.4_{1.2}$ $87.1_{0.2}$
-$1$ Tang et al. [2024] 4 Tang et al. [2024]* GPT-3 babbage $64.1_{3.9}$ $81.2_{3.0}$ $50.7_{4.1}$ $46.3_{7.8}$ $69.2_{7.9}$
+$1$ Tang et al. (2024) 4 Tang et al. (2024)* GPT-3 babbage $64.1_{3.9}$ $81.2_{3.0}$ $50.7_{4.1}$ $46.3_{7.8}$ $69.2_{7.9}$
+4 This work Gemma 1.1 2B IT $74.9_{3.8}$ $80.9_{3.6}$ $36.7_{2.2}$ $34.1_{9.3}$ $78.7_{1.9}$
 Ours 4 This work Gemma 1.1 2B IT $75.9_{3.5}$ $75.1_{0.5}$ $39.2_{3.7}$ $47.1_{6.0}$ $84.5_{1.0}$
@@
 (*) Note: For the results reported in Tang et al.
-[2024], they use GPT-3 babbage (now deprecated; we use GPT-3 babbage-002) as the in-context learner, and use the top 100 logprobs for contextual calibration (only top 5 are available now); while not directly comparable, we report their results for context.
+(2024), they use GPT-3 babbage (now deprecated; we use GPT-3 babbage-002) as the downstream in-context learner, and use the top 100 logprobs for contextual calibration (only top 5 are available now).
+While not directly comparable, we report their results for context.
 Results.
@@
 Our results at 64 shots are comparable to real data at 64 shots.
-Notably, our synthetic data at 64 shots improves over real data at 4 shots – which is roughly an upper bound on the performance of methods limited to generating 4 examples (e.g., Tang et al.
-[2024]).
+Notably, our synthetic data at 64 shots improves over real data at 4 shots – a rough upper bound on the performance of methods limited to generating 4 examples (e.g., Tang et al.
+(2024)).
 We also improve over results reported in Tang et al.
-[2024], although we note that there are differences in the experimental setup.
+(2024) – however as there are differences in the experimental setup, we also report the results of our re-implementation.
+[^4] Specifically, we use their best hyperparameters (from Appendix E, Table 9 of Tang et al.
+(2024)) and algorithm, but with our model, prompt, and evaluation setup.
+We evaluate with GPT-3 babbage-002 which has a 16K context window.
+We report results on AGNews, DBPedia, TREC, MIT-G, and MIT-D using the implementation of Zhao et al.
+(2021).
+Following the work of Tang et al.
+(2024), we enable contextual calibration (Zhao et al., 2021) for classification but not extraction tasks.
+Our evaluation setup is a best-effort reproduction of their setup, which is no longer possible to completely reproduce due to changes to OpenAI API access (see Table 2 caption).
+Due to cost, we follow prior work (Bertsch et al., 2024; Ratner et al., 2023; Lu et al., 2022; Zhao et al., 2021) and opt to subsample test sets to 250 test examples.
+We run 3 seeds of sampling of exemplars from synthetic/real data.
+Additionally, we present a limited set of results on Gemma 2 2/9/27B IT, studying the effect of model size on classification performance in §A.2.
 5.2 Fine-tuning
@@
 We use our approach to generate a large quantity of synthetic data for the purposes of fine-tuning 110M BERT-Base models.
-We consider 3 classification tasks used in prior work on private fine-tuning [Kurakin et al., 2024]), following the exact same evaluation procedure.
+We consider 3 classification tasks used in prior work on private fine-tuning (Kurakin et al., 2024)), following the exact same evaluation procedure.
 We omit comparison to prior private prediction work (e.g.
-[Tang et al., 2024]), as they only generate 4 examples which is insufficient for fine-tuning.
+(Tang et al., 2024)), as they only generate 4 examples which is insufficient for fine-tuning.
 Results.
@@
 Across various datasets and privacy levels, we generate between 2.5K (IMDB, $\varepsilon=1$) and 200K (Yelp, $\varepsilon=10$) examples for fine-tuning.
-Prior work generating fewer than $10$ examples using private prediction were unable to compare with private fine-tuning on these tasks at all.
+Prior work generating fewer than $10$ examples using private prediction were unable to compare with private fine-tuning on these tasks.
 While there remains a gap between the best fine-tuning and best private inference methods on downstream classification tasks, we achieve reasonable performance, even out-performing or matching the baseline of privately tuning all the parameters in the model reported in Kurakin et al.
-[2024].
+(2024).
 BERT Acc. (%)
@@
 Method Reported in Model $\infty$ $1$ $3$ $10$infty $1$ $3$ $10$infty $1$ $3$ $10$
-Real data [Kurakin et al., 2024] - $93.7_{0.1}$ $97.6_{0.1}$ $93.7_{0.1}$
-Fine-tune [Kurakin et al., 2024] LaMDA 8B $93.2_{0.2}$ $79.1_{1.7}$ $83.9_{0.6}$ $84.0_{0.7}$ $95.9_{0.1}$ $84.1_{0.3}$ $84.6_{0.1}$ $84.2_{0.3}$ $91.1_{0.1}$ $65.7_{2.9}$ $65.3_{2.7}$ $65.1_{5.3}$
+Real data (Kurakin et al., 2024) - $93.7_{0.1}$ $97.6_{0.1}$ $93.7_{0.1}$
+Fine-tune (Kurakin et al., 2024) LaMDA 8B $93.2_{0.2}$ $79.1_{1.7}$ $83.9_{0.6}$ $84.0_{0.7}$ $95.9_{0.1}$ $84.1_{0.3}$ $84.6_{0.1}$ $84.2_{0.3}$ $91.1_{0.1}$ $65.7_{2.9}$ $65.3_{2.7}$ $65.1_{5.3}$
 Prompt-tune $92.0_{0.1}$ $88.1_{0.4}$ $87.4_{0.2}$ $90.7_{0.2}$ $93.9_{0.1}$ $94.1_{0.1}$ $93.5_{0.1}$ $94.1_{0.1}$ $88.3_{0.3}$ $83.9_{0.8}$ $86.2_{0.2}$ $86.9_{0.1}$
@@
 We report mean and standard deviation over 3 runs of downstream fine-tuning and evaluation.
-We compare to results reported in [Kurakin et al., 2024] that fine-tunes a synthetic data generator with DP-SGD.
+We compare to results reported in (Kurakin et al., 2024) that fine-tunes a synthetic data generator with DP-SGD.
 We generate 2.5-200K examples with private prediction, which suffices to train reasonably performing models on.
@@
 We additionally consider the limited data regime.
-In Appendix A we present experiments on AGNews1K, a 1024-subsample of AGNews.
+In §A.1 we present experiments on AGNews1K, a 1024-subsample of AGNews.
 Our method, which employs parallel composition, is “pay-as-you-go”, i.e., we can put in a small amount of data to get out a small amount, while preserving quality.
@@
 This means that without sufficient data, all outputs will be low quality.
-Results in Table 5 demonstrate that our private prediction method generates more useful examples for in-context learning in this regime.
+Results in Table 5 demonstrate that our private prediction method generates more useful examples for in-context learning in this limited data regime.
 5.3 Structured data
@@
 Experiment setup.
-For JSON generation, we evaluate on a dataset of information about American movies scraped from Wikipedia [Rust, 2024].
+For JSON generation, we evaluate on a dataset of information about American movies scraped from Wikipedia (Rust, 2024).
 Entries contain fields such as title, year, cast, and extract (a short synopsis).
@@
 6 Discussion
-We believe that our significantly improved performance relative to Tang et al. [2024] is primarily attributable to two algorithmic innovations.
+We believe that our significantly improved performance relative to Tang et al. (2024) is primarily attributable to two algorithmic innovations.
 First, for each generated token, Tang et al.
-[2024] preserve the privacy of the entire distributions from which the token is sampled (by taking argmax), even though only the token itself is included in the synthetic data.
+(2024) preserve the privacy of the entire distributions from which the token is sampled (by taking argmax), even though only the token itself is included in the synthetic data.
 By contrast, our method uses a discrete choosing mechanism, the exponential mechanism.
@@
 In order to switch to the Gaussian mechanism using its standard $(\varepsilon,\delta)$-DP guarantee, and achieve comparable privacy guarantees we would would require $\sigma\approx 0.53$ to achieve a comparable privacy guarantee.
-(See Appendix D).
+(See §E).
 Better analyses of the Gaussian mechanism exist, but do not offer much help.
-Using the improved analysis in Balle and Wang [2018] to attain the same $\varepsilon$ would require $\sigma\approx 0.34$.
+Using the improved analysis in Balle and Wang (2018) to attain the same $\varepsilon$ would require $\sigma\approx 0.34$.
 Conducting the analysis so that both mechanisms have equivalent privacy loss under zCDP yields $\sigma=0.2$.
@@
 Secondly, Tang et al.
-[2024] generated each token using a different random sample of the sensitive prompts, which is computationally very expensive, as it prevents the use of KV cache-accelerated decoding, since the cache is invalidated upon every resample.
+(2024) generated each token using a different random sample of the sensitive prompts, which is computationally very expensive, as it prevents the use of KV cache-accelerated decoding, since the cache is invalidated upon every resample.
 While resampling less often would be more practical, Tang et al.
-[2024] noted that in this case the privacy amplification benefits of subsampling would not be adequately realized, and characterized this limitation as the “main weakness” of their approach.
+(2024) noted that in this case the privacy amplification benefits of subsampling would not be adequately realized, and characterized this limitation as the “main weakness” of their approach.
 Instead, our method generates each synthetic example using a fixed disjoint subset of the sensitive prompts, allowing us to leverage parallel composition in our analysis, and thus avoid this privacy versus computation tradeoff.
@@
 Limitations
-While our work demonstrates that private prediction is a practical technique for privately generating a large volume of high-quality synthetic data, there remains a small gap between our results and the results obtained from privately fine-tuning the parameters of the LLM.
+While our work demonstrates that private prediction is a practical technique for privately generating a large volume of high-quality synthetic data, there remains a gap between our results and the results obtained from privately fine-tuning the parameters of the LLM.
 Currently, private prediction methods pay a privacy cost for every generated token, while private fine-tuning methods do not.
-We view correcting this limitation as a very important open problem.
 Finally, any method for ensuring data privacy will inevitably entail some loss of data utility.
@@
   • Everyone contributed to discussing, interpreting, and iterating on experiment results as well as project management.
-Appendix A Private prediction beats fine-tuning in the limited data regime
+Appendix A Additional experiments
+A.1 Private prediction beats fine-tuning in the limited data regime
 We do LoRA fine-tuning with DP-SGD on AGNews1K, with the same setup that beats our method in the full data regime.
@@
 We evaluate performance on 4 and 16 shot in-context learning with GPT-3 babbage-002 (the same experimental setting as §5.1).
+Our private prediction approach outperforms the fine-tuning setup that does better in the full data regime.
 $\varepsilon$ Method Shots Model Acc. (%)
@@
 16 $68.1_{5.9}$
-Ours 4 Gemma 2B 1.1 IT $73.9_{8.3}$
+Ours 4 Gemma 1.1 2B IT $73.9_{8.3}$
 16 $80.1_{2.5}$
@@
 On the other hand fine-tuning does worse due to the stricter requirement that all future model outputs must be private.
-We run 16 shot since we only generate 38 examples/since 16 examples fills up the context length for LoRA.
+Evaluation setup is the same in §5.1, except here we run 16 instead of 64-shot, because 16 examples produced by the LoRA model fills up the entire context length of babbage-002.
+A.2 Effect of model size
+We report results on the effect of the data generator’s size on in-context classification performance on DBPedia.
+Our setup is the same as the experiments in §5.1, with the change that we use the Gemma 2 2/9/27B IT models to get more size variation in the same model family.
+This necessitated a slight change in the prompt (specifically, we append to the instruction: “No formatting or explanations.” For this evaluation, we see limited improvement due to scale.
+$\varepsilon$ Shots Model Acc. (%)
+1 4 Gemma 2 2B IT $75.7_{0.8}$
+64 $91.2_{0.0}$
+4 Gemma 2 9B IT $76.4_{1.2}$
+64 $92.4_{1.4}$
+4 Gemma 2 27B IT $76.9_{2.2}$
+64 $91.9_{1.9}$
+Table 6: Results on DBPedia classification. Evaluation setup is the same as in §5.1. We see limited improvement from the increase in model size.
 Appendix B Design choices
@@
 The baseline approach it to clip all logits to the interval $[-c,c]$ before aggregation and softmax – we refer to this as “fixed interval clipping”.
-Alternatively, we can clip to the range $[\max_{j}\{\mathbf{z}_{j}\}-2c,\max_{j}\{\mathbf{z}_{j}\}]$ and then translate to the interval $[-c,c]$ (Equation 1).
+Alternatively, we can clip to the range $[\max_{j}\{\mathbf{z}_{j}\}-2c,\max_{j}\{\mathbf{z}_{j}\}]$ and then translate to the interval $[-c,c]$ (Eq.
+1).
 In Figure 3 we plot the distortion as a consequence of clipping in terms of L1 error, and find that the latter approach allows us clip more than twice as aggressively, thus improving the privacy guarantee, without compromising utility.
+Appendix C Privacy checklist
+We follow privacy reporting guidelines (Ponomareva et al., 2023, Section 5.3.3) to state all information needed to get a complete picture of our privacy guarantee.
+  1. 1.
+    DP setting. This a central DP guarantee where the service provider is trusted to correctly implement the mechanism.
+  2. 2.
+    Instantiating the DP definition.
+    1. (a)
+      Data accesses covered. DP guarantees only apply to a single synthetic dataset generation run. We do not account for hyperparameter tuning.
+    2. (b)
+      Final mechanism output.
+The mechanism’s output is the full sequence of sampled tokens of the output synthetic dataset (all synthetic data instances), which can viewed as responses to a sequence of adaptive queries with different input text prefixes.
+    3. (c)
+      Unit of privacy. Example-level DP, where an example is the full sequence of tokens for one input sensitive example.
+    4. (d)
+      Adjacency definition for neighboring datasets. Add/remove one example.
+  3. 3.
+    Privacy accounting details.
+    1. (a)
+      Type of accounting used. zCDP based accounting.
+    2. (b)
+      Accounting assumptions.
+Our analysis assumes independent assignment of each prompt to a batch (Assumption 1).
+We can fulfill this with hashing (see §4), which corresponds to assigning each prompt to a batch uniformly at random.
+However, the implementation uses fixed size batches constructed via shuffling.
+    3. (c)
+      The formal DP statement. We report $(\varepsilon,\delta)$-DP for $\varepsilon=1,3,10$ and $\delta=(\texttt{training\_set\_size})^{-1}$.
+  4. 4.
+    Transparency and verifiability.
+We provide a full description of the algorithm and a complete privacy analysis.
+Our implementation is built off of the official open-source JAX library for Gemma: https://github.com/google-deepmind/gemma.
+At this time, the implementation is close-sourced.
+Discussion.
+The full output synthetic dataset of our mechanism indeed satisfies example-level add/remove DP.
+Below we discuss the more granular aspects of our privacy guarantee that are not captured in the above statement.
+First, as a consequence of parallel composition, we note that each individual synthetic example output only incurs privacy cost on the batch used to generate it.
+Hence, having the mechanism release a subset of the output will only incur privacy cost on a subset of our input dataset.
+Second, our current privacy guarantee holds with respect to the change of an entire context, which may consist of multiple examples.
+However, as we do not prescribe a way to group examples into contexts, we cannot a priori identify which examples belong to a single context.
 (a) Distribution of L1 error induced by fixed interval clipping.
@@
 We observe less error, at lower choices of $c$ when clipping with recentering (note the $x$-axis scales).
-Appendix C Proof of Theorem 1
+Appendix D Proof of Theorem 1
 Our proof of Theorem 1 is organized into sections.
-§C.1 provides basic definitions.
-§C.2 and §C.3 establish key results related to composition and sensitivity.
-§C.4 proves the privacy of simpler mechanisms that each account for a portion of the functionality of Algorithm 1.
-C.5 puts all the pieces together and completes the proof.
-C.1 Definitions
+§D.1 provides basic definitions.
+§D.2 and §D.3 establish key results related to composition and sensitivity.
+§D.4 proves the privacy of simpler mechanisms that each account for a portion of the functionality of Algorithm 1.
+§D.5 puts all the pieces together and completes the proof.
+D.1 Definitions
 In §4 we defined neighboring prompt datasets. We extend the definition to arbitrary sets.
@@
 Zero-concentrated differential privacy (zCDP) is a relaxation of $\varepsilon$-differential privacy.
-Definition 4 (Bun and Steinke [2016]).
+Definition 4 (Bun and Steinke (2016)).
 A mechanism $M$ satisfies $\rho$-zCDP if
@@
 for all $\alpha>1$ and neighboring datasets $D,D^{\prime}\in\mathcal{D}$, where $D_{\alpha}(P\parallel Q)$ is Rényi divergence of order $\alpha$ betweeen distributions $P$ and $Q$.
-C.2 Composition
+D.2 Composition
 Zero-concentrated differential privacy obeys a simple sequential composition rule.
@@
 If mechanisms $M_{1}$ and $M_{2}$ satisfy $\rho_{1}$-zCDP and $\rho_{2}$-zCDP, respectively, then the sequential composition of $M_{1}$ and $M_{2}$ satisfies $(\rho_{1}+\rho_{2})$-zCDP.
-Parallel composition is a well-known technique in differential privacy that is useful for establishing privacy guarantees in scenarios where a mechanism is independently applied to disjoint subsets of a dataset.
-Many versions of parallel composition require that the subsets are chosen in a fully data-independent manner.
+Parallel composition is a well-known technique in differential privacy that is useful for establishing privacy guarantees in scenarios where a mechanism is independently applied to disjoint batches of a dataset.
+Many versions of parallel composition require that the batches are chosen in a fully data-independent manner.
 We show that the same result holds under a weaker assumption.
 Lemma 2.
-Let $k$ be a positive integer. Let $f$ be a function that maps prompts into $[k]$. For any dataset of prompts $D$ and $i\in[k]$ let
+Let $k>1$ be the number of batches.
+Let $f$ be a batch assignment function that satisfies Assumption 1.
+In other words, let $f$ be a function that map each prompt to and element of $[k]$.
+For any dataset of prompts $D$ and $i\in[k]$ let
 $D_{i}=\{\mathbf{p}\in D:f(\mathbf{p})=i\}.$
-Let $M$ be a mechanism that satisfies $\rho$-zCDP. If $\widehat{M}$ is the mechanism defined by
+be the $i$th batch of prompts. Let $M$ be a mechanism that satisfies $\rho$-zCDP. If $\widehat{M}$ is the mechanism defined by
 $\widehat{M}(D)=(M(D_{1}),\ldots,M(D_{k}))$
@@
 We have for all $\alpha>1$
-$D_{\alpha}(\widehat{M}(D)\parallel\widehat{M}(D^{\prime}))$ $=\sum_{i=1}^{k}D_{\alpha}(M(D_{i})\parallel M(D^{\prime}_{i}))$
-$=D_{\alpha}(M(D_{j})\parallel M(D^{\prime}_{j}))$
+$D_{\alpha}(\widehat{M}(D)\|\widehat{M}(D^{\prime}))$ $=\sum_{i=1}^{k}D_{\alpha}(M(D_{i})\|M(D^{\prime}_{i}))$
+$=D_{\alpha}(M(D_{j})\|M(D^{\prime}_{j}))$
 $\leq\rho\alpha\qed$
-C.3 Sensitivity analysis
+D.3 Sensitivity analysis
 In this we compute the sensitivity of several functions used in Algorithm 1. Each function depends on a set of logit vectors. Recall that a logit vector is an element of $\mathbb{R}^{v}$. Let
@@
 where $\operatorname{clip}_{c}(\cdot)$ was defined in Eq. (1). Also recall the distance function defined in Eq. (2):
-$d(Z,\mathbf{z})=\left\lVert\frac{1}{s}\sum_{\mathbf{z}^{\prime}\in Z}p_{\mathbf{z}^{\prime}}-p_{\mathbf{z}}\right\rVert_{1}$
+$\mathrm{d}(Z,\mathbf{z})=\left\lVert\frac{1}{s}\sum_{\mathbf{z}^{\prime}\in Z}p_{\mathbf{z}^{\prime}}-p_{\mathbf{z}}\right\rVert_{1}$
 where $p_{\mathbf{z}}=\operatorname{softmax}(\mathbf{z})$.
 Lemma 3.
-The function $\ell$ has sensitivity $\frac{c}{s}$, and for all $\mathbf{z}\in\mathbb{R}^{v}$, the function $d(\cdot,\mathbf{z})$ has sensitivity $\frac{1}{s}$.
+The function $\ell$ has sensitivity $\frac{c}{s}$, and for all $\mathbf{z}\in\mathbb{R}^{v}$, the function $\mathrm{d}(\cdot,\mathbf{z})$ has sensitivity $\frac{1}{s}$.
 Proof.
@@
 We also have
-$\left|d(Z,\mathbf{z})-d(D^{\prime},\mathbf{z})\right|$
+$\left|\mathrm{d}(Z,\mathbf{z})-d(D^{\prime},\mathbf{z})\right|$
 $=$ $\left|\left\lVert\frac{1}{s}\sum_{\mathbf{z}^{\prime}\in Z}p_{\mathbf{z}^{\prime}}-p_{\mathbf{z}}\right\rVert_{1}-\left\lVert\frac{1}{s}\sum_{\mathbf{z}^{\prime}\in Z^{\prime}}p_{\mathbf{z}^{\prime}}-p_{\mathbf{z}}\right\rVert_{1}\right|$
@@
 where we used the reverse triangle inequality. ∎
-C.4 Constituent mechanisms
+D.4 Constituent mechanisms
 In this section we prove privacy guarantees for several simpler mechanisms that we will later compose together to show that Algorithm 1 is private.
@@
 Consider a function $f:\mathcal{D}\rightarrow\mathbb{R}^{v}$ with sensitivity $\Delta$.
-By an analysis of the exponential mechanism due to Cesar and Rogers [2021], [^6] See also Rogers and Steinke [2021].
+By an analysis of the exponential mechanism due to Cesar and Rogers (2021), [^6] See also Rogers and Steinke (2021).
 choosing a token according to the distribution $\operatorname{softmax}(\frac{\varepsilon}{2\Delta})$ satisfies $\frac{1}{8}\varepsilon^{2}$-zCDP.
@@
 7:     $\mathbf{z}_{\operatorname{public}}\leftarrow\operatorname{logits}(\mathbf{p}_{\operatorname{public}}\mathbf{x})$
-8:     $\hat{d}\leftarrow d(Z,\mathbf{z}_{\operatorname{public}})+\textrm{Laplace}(2\sigma)$
-9:     if $\hat{q}>\hat{\theta}$ then
+8:     $\hat{d}\leftarrow\mathrm{d}(Z,\mathbf{z}_{\operatorname{public}})+\textrm{Laplace}(2\sigma)$
+9:     if $\hat{d}\geq\hat{\theta}$ then
 10:         Break
@@
 Proof.
-Observe that mechanism $M$ is an instance of the AboveThrehold mechanism [Dwork et al., 2009], which accepts a private dataset, a threshold, and a sequence of queries as input.
+Observe that mechanism $M$ is an instance of the AboveThrehold mechanism (Dwork et al., 2009), which accepts a private dataset, a threshold, and a sequence of queries as input.
 In each iteration, the AboveThreshold mechanism applies the next query in the sequence to the dataset and compares it to a noisy threshold, and returns the index of the first query that exceeds the threshold.
@@
 Thus by the analysis due to Dwork et al.
-[2009], mechanism $M$ satisfies $\frac{2}{s\sigma}$-differential privacy, which by Bun and Steinke [2016] implies that mechanism $M$ satisfies $\frac{2}{(s\sigma)^{2}}$-zCDP.
+(2009), mechanism $M$ satisfies $\frac{2}{s\sigma}$-differential privacy, which by Bun and Steinke (2016) implies that mechanism $M$ satisfies $\frac{2}{(s\sigma)^{2}}$-zCDP.
 ∎
-C.5 Putting it all together
+D.5 Putting it all together
 Consider a sequence of iterations of the inner loop of Algorithm 1 in which the value of $t$ (the private token counter) is constant.
@@
 It remains to convert this zCDP guarantee to an $(\varepsilon,\delta)$-differential privacy guarantee, which we do two different ways using two existing results: Corollary 13 due to Canonne et al.
-[2020] and Lemma 3.5 due to Bun and Steinke [2016].
-Appendix D Privacy-equivalent Gaussian noise
+(2020) and Lemma 3.5 due to Bun and Steinke (2016).
+Appendix E Privacy-equivalent Gaussian noise
 Given the average token distribution $\bar{\mathbf{p}}$ in a batch, Tang et al.
-[2024] protect the privacy of $\bar{\mathbf{p}}$ by using the Gaussian mechanism, which achieves $(\varepsilon,\delta)$-differential privacy with $\varepsilon=\frac{\sqrt{2\log(1.25/\delta)}}{s\sigma}$, where $s$ is the batch size and $\sigma$ is the standard deviation of the noise added to each probability in $\bar{\mathbf{p}}$.
+(2024) protect the privacy of $\bar{\mathbf{p}}$ by using the Gaussian mechanism, which achieves $(\varepsilon,\delta)$-differential privacy with $\varepsilon=\frac{\sqrt{2\log(1.25/\delta)}}{s\sigma}$, where $s$ is the batch size and $\sigma$ is the standard deviation of the noise added to each probability in $\bar{\mathbf{p}}$.
 On the other hand, we use the exponential mechanism to protect the privacy of a sample drawn from $\bar{\mathbf{p}}$, which achieves $\varepsilon$-differential privacy with $\varepsilon=\frac{2c}{s\tau}$, where $c$ is the maximum absolute value of any log-probability in the batch and $\tau$ is the sampling temperature.
@@
 Plugging in the above parameters yields $\sigma\approx 0.53$.
-The analysis in Theorem 8 of Balle and Wang [2018] does not admit a closed-form solution. Instead, we binary search for a solution to:
+The analysis in Theorem 8 of Balle and Wang (2018) does not admit a closed-form solution. Instead, we binary search for a solution to:
 $\Phi\left(\frac{\Delta}{2\sigma}-\frac{\varepsilon\sigma}{\Delta}\right)-\exp(\varepsilon)\Phi\left(-\frac{\Delta}{2\sigma}-\frac{\varepsilon\sigma}{\Delta}\right)\leq\delta$
@@
 This procedure yields $\sigma\approx 0.34$.
-Finally, equating the zCDP loss for the exponential mechanism given by $\frac{\varepsilon^{2}}{8}=\frac{c^{2}}{2s^{2}\tau^{2}}$ (Cesar and Rogers [2021]) to that of the Gaussian mechanism given by $\frac{1}{2s^{2}\sigma^{2}}$ (Bun and Steinke [2016]), yields $\sigma=0.2$.
-Appendix E Experiment details
-E.1 Hyperparameter tuning
+Finally, equating the zCDP loss for the exponential mechanism given by $\frac{\varepsilon^{2}}{8}=\frac{c^{2}}{2s^{2}\tau^{2}}$ (Cesar and Rogers (2021)) to that of the Gaussian mechanism given by $\frac{1}{2s^{2}\sigma^{2}}$ (Bun and Steinke (2016)), yields $\sigma=0.2$.
+Appendix F Experiment details
+F.1 Hyperparameter tuning
 There are a significant amount of hyperparameters associated with our approach.
@@
 So for $\tau=2.25$ settings only, we follow Tang et al.
-[2024] and reduce the vocabulary to the public prediction’s top 1024.
+(2024) and reduce the vocabulary to the public prediction’s top 1024.
 We emphasize that (1) we do not do this for any of the other settings of $\tau$, and (2) use a larger value than prior work (they use top 100).
@@
 Table 7: Values for hyperparameters explored in this work.
-E.2 Prompts used
+F.2 Prompts used
 We report the prompts used for our experiments.
@@
 Figure 12: Public generation prompt for WikiMoviesJSON.
-Appendix F Artifacts
+Appendix G Artifacts
 Tables 1(a) and 1(b) list all artifacts we use in this work.
 AGNews, TREC, DBPedia, MIT-G, MIT-D, IMDB, and Yelp are all standard academic datasets permissible for research use; we cite their original publications when introduced.
-WikiMoviesJSON is scraped from Wikipedia data, courtesy of [Rust, 2024]; their work is covered by an MIT license.
+WikiMoviesJSON is scraped from Wikipedia data, courtesy of (Rust, 2024); their work is covered by an MIT license.
 Wikipedia content is licensed under the Creative Commons Attribution-ShareAlike 4.0 International License (CC BY-SA) and the GNU Free Documentation License (GFDL).
-We use open-source models BERT-Base, released by [Turc et al., 2019], and Gemma.
+We use open-source models BERT-Base, released by (Turc et al., 2019), and Gemma.
 Our use of Gemma for academic purposes is in accordance of the Gemma terms of use: https://ai.google.dev/gemma/terms.
@@
 LaMDA 8B is not publically available, but we received sufficient authorization to use it for the academic purposes of this paper.
-Appendix G Compute budget
-All experiments for synthetic data generation run on Gemma 2B 1.1 IT.
+Appendix H Compute budget
+Our main experiments for synthetic data generation run on Gemma 1.1 2B IT.
 A run of synthetic data generation takes between 8-48 accelerator hours.